A Bibliography of Publications about the C# Programming Language

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

01 February 2024 Version 1.59

Title word cross-reference

$49.99 [Nau08a]. $54.99 [Nau08b]. ^{♯} [KR18].

.NET [Fox04, Hol04, Nau08b, SKS08, AHG^{+}04, ADM02, Ano02, Ano03b, Bal03, Bar03, BHW05, BS03, Bro09, Bui02, BW02, CSH03, CAGH02, FM02, FLF08, FJ03, GS05, Gri12b, Gri02, LN04, LRT^{+}18, LH04, LM06, LL03a, LL03b, Mac08, May08, Mic09, NEG^{+}10, Nil06, Oni03b, PH03, Per06, PW01, RW02, SG04, SJNH02, Shu05, SSRW02, Sub05, Tro01, Tro03, Tro05, UKH^{+}04, Vic04, WAH03, YD04, Gro22, Tro10].

0-521-68761-6 [FV09].

1 [Nau08b]. 1-59059-955-1 [Nau08b]. 1.1 [Mac05a]. 10 [Alb22]. ’18 [ACM18].

2 [BRE06a, BRE06b]. 2.0 [HSH04, Mic07, Sch06a, Tro05]. 2001 [ACM01]. 2003 [Fox04, Hol04, Int03, Int05a]. 2005 [Bak06, Bal06, Dav05, DS06, FH06, Fox06, Hof06, JM06, Ken05, Lho06, Lib05b, Mar05, Mur06, Pel05, Pel06a, Sha06b, Sha06a, Tro05, W^{+}05, Wat06, W^{+}06]. 2006 [Int06]. 2008 [DD09, Fox08, Mac08, Nau08a, SD08]. 2010 [DD11, JF10, Mac10, Tro10, Wat10b]. 23272 [Int05b]. 2nd [Nau08b].

3.0 [AAD07, AA08a, Ano05b, HT07, Hum06, LX08, May08, Mic09]. 3.5 [Mac08, May08, Mic09, Nau08b]. 334 [C#02, ECM01, ECM02a, ECM05, ECM06, ECM02b].

4 [ERH10, NEG^{+}10]. 4.0 [AA10a, AA10b, GAL10, Mag10, Mic10, Sch10, Tro10, Wat10a].

5 [Nau08a]. 5.0 [Gri12a, Gri12b].

6 [Gro22].

70-306 [Ano02]. 70-310 [Ano03b]. 70-316 [Ano02]. 70-320 [Ano03b].

8 [Ano17].

95 [CSH03]. 978 [Ano17, Nau08a]. 978-1-59059-965-5 [Nau08a]. 978-1-59327-759-8 [Ano17].

absolute [Bak06]. Abstract [McC04c, SSK13]. Abstraction [BS04]. abstractions [BCF04, SPP^{+}10]. access [Hum06, Wil03]. Acid [RIS^{+}17]. Acronyms [Bar01]. activity [CKL^{+}13]. Ada [CSH03, KPPÉR06]. adaption [AK09]. adaptive [SVY09]. ADO.NET [HSF03, VB02, Wil03]. adoption [Ano03a]. Advanced [Kim02]. again [STW15]. Agent [CDP16]. Agent-Based [CDP16]. aggressive [MGM^{+}06]. AI [Lut03]. algebraic [KR18]. alternative [Ano05a]. Alternatives [Swa01]. Alto [ACM01]. Alzheimer [CDP16]. Analysemodell [BRE06b]. Analysis [Doy10, Ano05b, BRE06b, Doy08, FV09, GBJ^{+}10, GBJ^{+}12, GBJ^{+}13, Gro22, McS10, SD12, Ano05b]. Anchor [MSK09]. Anders [Bar01]. Annotated [AHG^{+}04, JPS07, MR04]. annotations [Kic04, PVR^{+}20]. Anonymity [Bar01]. apples [BNK^{+}07]. application [CKL^{+}13, Dun03]. Applications [Gri12b, Ano02, Ano04a, CAGH02, CMLC06, LH04, PH03, SPP^{+}10]. Applied [DR12, SSRW02, Wor04]. Applying [Lut03, Nil06]. Approach [JM06, JF10]. Apress [Nau08a, Nau08b]. Architecture [Wan11, Bal03]. Aren’t [BHP^{+}01]. Arithmetic [ACM18]. ASP.NET [Mac05b, BH02, ERH10, GKM^{+}02, GKM^{+}03, Hof03, HSH04, KMM^{+}02, LH02, LH03, Mac05a, Oni03b, Oni03a, Rei02, Ruv04, UKH^{+}04]. assembly [TCH10]. Assessing [Mey02]. asymmetric [CBGM12]. asynchronous [SK17]. atomic [ZHU^{+}10]. Atomicity [JLP^{+}14, AHM09]. Attributes [Kic04]. Audits [Ano05b]. Ausdrücke [SKS08]. automata [ZWZ^{+}14]. Automated [HP09, YH10]. automatic [SB07]. Award [Eri01].

Bad [BHP^{+}01, BNK^{+}07]. Base [AHG^{+}04, Bal06]. Based [CDP16, RIS^{+}17, AZLY18, Ano02, FV09]. Basic [Ano02, Ano03b, Gro22, PH03, SG04, Vic04, MSK09, Ano04a, BD05, Oni03b, Rus08]. Batteries [RIS^{+}17]. battlefield [WT10]. been [Hun03]. beginner [Bak06, Buo03, Buo05, Sch01]. Beginning [BP04, BW02, CAGH02, CAGH04, FL02, GKM^{+}02, GKM^{+}03, HRA^{+}05, KMM^{+}02, Mac05a, Pen05, UKH^{+}04, WEG^{+}02, W^{+}05, Wat06, W^{+}06, Wat10b]. best [BD05, VB02]. beta [H^{+}00]. better [ABE^{+}02]. between [MP18]. bible [FPB^{+}02]. Bibliography [Bee06]. bis [BRE06b]. Black [Tel06, Tel02, Wag02]. Blame [STW15]. blocks [ZHU^{+}10]. Bluetooth [SG01]. bodies [MP18]. Body [GBJ^{+}13, GBJ^{+}10, GBJ^{+}12]. Book [Ano17, Cir01, FV09, Tel06, Tel02, Wag02]. Books [Ano01]. Bookshelf [Gri02, Lut03, SG01, Wil01]. both [Ano04b]. Brandon [Ano17]. Bridges [Ano04a]. brings [Ano05b]. Build [Ruv04, Pel05]. Building [Ano04a, Gri12b, Lut03, PH03, Ano03a]. built [Ano04a]. Business [Lho06]. Bytecode [LN04]. Bytecode-to-.NET [LN04].

C [Cir01, Ano05b, Bru05, G^{+}01, GK03, Hum03, KR18, Kic04, MP18, Nod04, PW01, Ste02a, Ste02b]. C# [Ano02, Ano03b, Bak06, BD05, Bal06, Bra06, DD11, FH06, Fox04, Fox06, Fox08, Hof06, Hol04, JM06, JF10, Ken05, Lho06, Lib05b, Mar05, Mös06, Nau08a, Nau08b, Pel05, Pel06a, Sch06b, SJNH02, Sha06b, Sha06a, SKS08, Tro10, UKH^{+}04, W^{+}05, W^{+}06, Wat10b, PVR^{+}20, ADM01, ADM02, AAD07, AA08a, AA10a, AA10b, Alb22, ABE^{+}02, Ano01, Ano03a, Ano04a, Ano04b, Ano05a, Ano05b, AW02, Arc02, AWvS^{+}02, Bak02, BGM06, Bar01, BP04, Bar06, Bay07, Bay08, Bee06, BCF04, Bey01, BMT07, BH04a, Bis04, BHW05, Bis07, Bo02, BHP^{+}01, BS04, BFGS05, Bro09, Bru05, Buo03, Buo05, BW02, CAGH02, CAGH04, CG05, CDP16, CDH^{+}03, Coo03, Cro01, C#02, Dav05, DS06, Dei02, D^{+}03, DD06a, DD06b, DD09, DS12, DR12, Doy08, Doy10, DLW05, DAN02b]. C# [DAN02a, DAN03, DLE06, Dun03, ECM01, ECM02a, ECM05, ECM06, ECM02b, Ead01, ERH10, FPB^{+}02, FHCM^{+}02, FL02, Fru10, G^{+}01, GC15, GS05, GK03, GKM^{+}02, GKM^{+}03, GAL10, Gri12a, Gri12b, Gri02, GBJ^{+}10, GBJ^{+}12, GBJ^{+}13, Gro22, GN01, GWH05, HP04, H^{+}00, HWG04, HWG06, HTWG08, HTWG11, Her02, HT06, HT07, HRA^{+}05, Hum06, Hum03, Hun03, Int03, Int06, JPS07, JJOA07, Jon02, Jon04, KPPÉR06, KMM^{+}02, Kic04, Kim02, Lau17, Lew06, Lib01, Lib02a, Lib02b, Lib03, Lib05a, LM06, LM08, LX08, Lip01, Lut03, LL03a, LL03b, Mac04, MJ03, Mac05a, Mac05b, Mac08, Mac10, Mag10, Mac06, May08, McC03a, McC03b, McC03c, McC03d, McC03e, McC04a, McC04b, McC04c, McC04d, McC04e, McC05, Met04, Mey02, Mic07, Mic09, Mic10, Moj02, MP18, Mös04, ML04, Mur06]. C# [NEG^{+}06, NEG^{+}10, Nil06, Nod04, Oni03a, Pel06b, Pen05, PS17, Per06, PSN^{+}02, Pet02, PW01, Reg02a, RHSvS02, RW02, Ric06, Ric10, R^{+}02, Rob02, RAC^{+}03, RNG^{+}04, Ruv04, SK17, SK08, Sax03, Sch01, Sch06a, Sch10, Sel04, SH04, Shu05, SSK13, Ske10, SSRW02, SD08, SG01, Stä05, SG08, SG10, Ste02a, Ste02b, Sub05, Tag03, TH04, THG^{+}02, Tel02, Tel06, Tho02, TFRS03, Tro01, Tro03, Tro05, VB02, Wag02, Wag05, Wag10, WEG^{+}02, Wat06, Wat10a, WLR05, WdC01, WNG03, Wil00, Win04, Wor04, YD04, dCP05, Ano17, FV09]. C#-2005-Codebook [Bay07]. C/C [Cir01]. Cache [CS06]. Cache-conscious [CS06]. Calif [ACM01]. Call [ZWZ^{+}14]. Callbacks [Hum03]. calling [ZWZ^{+}14]. Campbell [FV09]. Can [Reg02a]. Capabilities [DS12]. Case [BS04, HNTL12, KPPÉR06]. Centre [ACM18]. certifying [MSLL07]. Chameleon [SVY09]. change [GC15]. chaos [GBJ^{+}10, GBJ^{+}12, GBJ^{+}13, GBJ^{+}13]. checking [XJC09]. checks [SB07]. choose [Ano04b]. Class [AHG^{+}04, Bal06, BHP^{+}01, AK09, CDH^{+}03, Fos05]. Classes [Fos05, McC03e, McC04c, PS17, CDH^{+}03, Fek08, JJOA07, Tit13, TWBS09]. CLR [Ric06, Ric10]. coallocation [CS06]. Code [Ano05b, BHP^{+}01, NBV14, BP04, FLF08, GBJ^{+}13, PVR^{+}20, SSK13, GBJ^{+}10, GBJ^{+}12]. Codebook [Bay07, Bay08, Bay07, Bay08]. coding [CDH^{+}03]. coercion [STW15]. Colin [FV09]. collaborative [RRY10]. Collection [GMN21, KPHV11, PZM^{+}10]. Collections [McC04b, SVY09]. collectors [HP09, KPS09, MSLL07, PPS08]. college [Gou08]. column [Hun03]. COM [Bey01]. commerce [Hof03]. commercial [GC15]. Common [BS03, Int05a, Int05b, Cro01, Hun03, MR04, WLR05]. Compact [YD04]. companies [Ano04a, Ano04b]. Comparing [KPPÉR06]. Comparison [CDP16]. Competitor [Win04]. Compilation [JLP^{+}14, CHP^{+}08, JK13]. compiler [CMLC06, HP04, KSD12, MGM^{+}06]. compilers [CHP^{+}08]. complete [Bis04, Fox06, Fox08, Her02, Sch06a, Sch10, Tag03]. completeness [KBPS17]. completion [PGBG12]. Complex [GD07, Gou08]. Components [Rei02, Ano03b, PH03]. Computer [Bar01, GD07]. Computing [GK03]. concept [MSK09]. concepts [BP04, Ses08]. concisely [BH04a]. concurrency [BCF04]. concurrent [JPS^{+}08, KPS09, PPS08]. Conference [ACM01, ACM18]. CoNGA [ACM18]. conscious [CS06]. Construction [LN04]. contract [XJC09]. Contributing [Cir01]. controls [Mac06]. Convention [ACM18]. convergence [KSD12]. Conversion [Moj02]. convex [MP18]. cookbook [HT06, HT07, TH04]. Core [Per06, Wag02]. Correction [BHP^{+}01]. Course [Wan11, RHSvS02, SJNH02]. creating [ABE^{+}02]. Creation [CDP16]. Crosscut [Kic04]. CS1 [ABE^{+}02, Reg02a]. CS2 [ABE^{+}02, Reg02a]. CSharp [BRE06a, BRE06b]. CUP [FV09]. Curb [JJOA07]. curriculum [Gou08]. custom [Mac06]. cycle [CKL^{+}13].

dager [Jon02]. Data [Lut03, CS06, Hum06, KR18, MJ03, McS10, MSK09, SB07, Wil03]. Database [Tho02]. Databases [Wat06, HRA^{+}05, KMM^{+}02]. days [Jon04].

DBA [Lut03]. deadlock [SK17]. Debugging [ZHU^{+}10]. deep [Ano05b]. definition [BFGS05]. Definitive [Alb22]. delayed [FX07]. Delegate [Lip01]. Demystified [Ken05]. Department [BHP^{+}01]. dependencies [TWBS09]. deployment [RW02]. Depth [Ske10]. descent [Sax03]. described [Hun03]. Design [Doy10, Lut03, Met04, RIS^{+}17, BH02, Bis07, CDH^{+}03, Coo03, Doy08, Hof03, Lew06, Nil06, RW02, CMLC06]. Desktop [Gri12b, DAN02a, DAN03]. detection [SK17]. develop [Fek08]. developer [DB04, Lib05b, Swe06]. developers [Ses08, BD05]. Developing [Ano03b, BH04b, Ano02, SSRW02]. Development [RIS^{+}17, Bak02, GD07, Gou08, RW02]. DHTML [BHP^{+}01]. diagram [CKL^{+}13]. Dijkstra [SWS^{+}13]. dimensional [MP18]. directed [PGBG12]. Disasters [Lut03]. Discussion [G^{+}01]. Disease [CDP16]. distance [MP18]. distributed [RRY10]. DITTO [SB07]. do [Ano04b]. Dobb [Eri01]. domain [Nil06]. domain-driven [Nil06]. down [ABE^{+}02]. Dr. [Eri01]. driven [Nil06]. Dumb [BHP^{+}01]. dummies [Dav05, DS06, SD08]. Dutch [R^{+}02, WNG03, FHCM^{+}02, GN01, PSN^{+}02, RHSvS02, SJNH02, WdC01]. Dynamic [Bar03, DLE06].

e-commerce [Hof03]. easy [Ano05a, Bak02]. easy-to-use [Ano05a]. ECMA [C#02]. ECMA-334 [C#02, ECM01, ECM02a, ECM05, ECM06, ECM02b]. ed [Nau08b]. Edition [Bak06, Bar06, Pel05, Pel06a, Pel06b]. Editors [Cir01]. effect [GC15]. Effective [Wag05, Wag10, CDH^{+}03]. efficient [NBV14]. einsteigen [FH06]. Einstieg [BRE06a]. elements [BGM06]. embedded [SD12]. empirical [PVR^{+}20]. Engine [GBJ^{+}13]. enterprise [RW02]. Entity [Ler10]. entrance [BRE06a]. environment [CSH03, Fos05, WAH03]. environment-passing [Fos05]. errors [BNK^{+}07]. Essential [BS03, Mic07, Mic09, Mic10, Oni03b, Oni03a]. Essentials [ADM01, ADM02]. Establishing [FX07]. étape [Sha06a]. EU [CDP16]. Evaluation [DS12, DR12, Wan11, PVR^{+}20, SA09]. Examining [McC03b]. examples [Nil06, Oni03b, Oni03a, VB02]. Exams [Ano02, Ano03b]. Excellence [Eri01]. Executables [BHP^{+}01]. exotic [GS05]. Experience [BHW05, BFL^{+}11, WT10]. expert [Mac05b, Lho06]. Exploiting [BS04]. Exploring [JK13]. Express [Bak06, Bar06, Pel05, Pel06b, Pel06a]. expressions [PGBG12]. Expressive [HS08]. Expressivity [KR18]. Extended [AHG^{+}04]. extending [Mag10]. extensible [McS10]. extensions [BMT07]. Extensive [Wan11].

F# [Shu05]. family [Kic04]. fast [Bak02, KCP^{+}17]. faster [TWBS09]. Fault [RV13, GC15]. fault-proneness [GC15]. financial [Wor04]. First [Ste02b, Fos05, HSF03, HSH04, STW15, SG08, SG10]. fixing [LTZ14]. flexible [DLE06]. Flint [LTZ14]. form [JJOA07]. Formal [Stä05]. formalizing [BMT07]. forms [Mac06, Sel04, SG04]. Fortify [Ano05b]. forum [Reg02b]. Foundation [Nau08b, Mac08]. foundations [Lew06]. FPGA [LRT^{+}18]. fractal [Gro22]. fragmentation [PZM^{+}10]. fragmentation-tolerant [PZM^{+}10]. Framework [AHG^{+}04, ADM02, LL03a, LL03b, May08, Mic09, YD04, MSLL07, PW01, Gri12b, Ler10]. free [KPS09]. French [Sha06a]. freshman [GD07]. functions [Tit13]. fusion [KBPS17]. future [Hum06].

G [KR18]. Game [Wan11, Buo03, Buo05, GD07, Gou08, Pen05, Swe06]. Garbage [GMN21, HP09, KPHV11, KPS09, MSLL07, PPS08, PZM^{+}10]. Garden [MSK09]. GCC [BHP^{+}01]. GDI [Cha04]. general [AZLY18, MSLL07]. Generalized [Hum03, KR18, WT10]. Generating [HBM^{+}06, SSK13]. Generation [ACM18, NBV14]. Generative [DLW05]. generators [Lau17]. Generic [GK03, TWBS09]. Generics [MPO08, SD12]. German [BRE06a, BRE06b, Bay07, Bay08]. Get [LM06]. GNU [GDT^{+}02]. Goes [Kic04, Ano04b]. gotchas [Sub05]. Grande [ACM01]. Grande/ISCOPE [ACM01]. Graphics [Cha04]. Gray [Ano17]. Groovy [AK09]. GUESSTIMATE [RRY10]. GUI [BH04b]. Guide [Ken05, Lut03, Bis04, Buo03, Buo05, MA03, Sch01, dCP05]. guidelines [BD05]. gut [SKS08].

handbook [CDH^{+}03, Mac04, MJ03, TFRS03]. Hardware [GMN21, AHM09, CBGM12]. Harmonizing [Tit13]. Haskell [XJC09]. Hastlayer [LRT^{+}18]. Head [SG08, SG10, Ano03a, Ano04b]. Hello [McC03b]. Hi [KSD12]. Hi-Lite [KSD12]. hierarchies [AK09]. High [LRT^{+}18, Lut03, BFGS05, G^{+}01]. High-level [LRT^{+}18, BFGS05]. High-Tech [Lut03]. higher [MPO08, SWS^{+}13]. higher-order [SWS^{+}13]. highly [SPP^{+}10]. hindered [Ano03a]. Hogging [Bar01]. hook [Kic04]. hot [CS06]. hours [FHCM^{+}02, Fox04, Fox06, Fox08]. how-to [Wat10a]. Huge [BHP^{+}01]. hyper [Gro22]. hyper-fractal [Gro22].

Ideas [BHP^{+}01]. idempotence [RV13]. IEC [Int03, Int05a, Int05b, Int06]. IEEE [LL03b]. III [Tit13]. Immersive [Lut03]. immutability [GPP^{+}12]. Implementation [LRT^{+}18, PS17, Bal03, CMLC06, Dun03]. Implementations [LRT^{+}18]. improve [Sub05, Wag05, Wag10]. incremental [KPS09]. incrementalization [SB07]. Independent [LN04]. Industry [Bar01, Reg02b]. Inferable [TCH10]. Information [Int03, Int05a, Int05b, Int06, LN04]. Infrastructure [Int05a, Int05b, Bro09, MR04, WLR05]. inleiding [GN01, WdC01]. inspiration [Dun03]. instruction [YH10]. Integrated [GMN21, McS10]. integrating [BHW05, SPP^{+}10]. Intel [BHP^{+}01]. Intelligence [Lut03]. intelligent [Ano05b]. interactive [MSK09]. Interfaces [McC04e, WT10]. Internet [Bar01, RIS^{+}17, Wil03]. interpreter [Fos05]. Interruptible [LKM06]. Introduction [McC03c, Bro09, D^{+}03, GN01, GWH05, WdC01]. invariant [SB07]. invariants [FX07]. ISAs [HNTL12]. ISB [FV09]. ISB-10 [FV09]. ISBN [Ano17, Nau08a, Nau08b]. ISCOPE [ACM01]. ISO [Int03, Int05a, Int05b, Int06]. ISO/IEC [Int03, Int05a, Int05b, Int06]. iterators [LKM06]. ivory [Reg02b].

J# [GS05]. Jacky [FV09]. Java [ACM01, SKS08, Ano03a, Ano04a, Ano04b, Ano05a, Ano05b, Bar01, Bar03, BHW05, BHP^{+}01, BS04, Bru05, DLE06, Ead01, Fek08, G^{+}01, GK03, Hun03, KPPÉR06, Kic04, LN04, Lut03, PVR^{+}20, Reg02a, SB07, SPP^{+}10, Swa01, Win04]. Java/C# [BS04]. JavaGI [WT10]. jetzt [Pel06b]. JIT [JK13]. Join [Rus08]. Jonathan [FV09]. JSP [Ano05b]. Julia [Lau17]. June [ACM01]. Just [Bar01].

keynote [McK16]. kick [Hol04]. Kids [Sch06b]. Killer [Bar01, Hun03]. kind [MPO08]. Kit [Bar06, Ano02, Ano03b, Fox06, Fox08]. kompakter [Mös06]. kurz [SKS08].

Language [Bee06, BS03, C#02, DAN02b, ECM01, ECM02a, ECM05, ECM06, ECM02b, HTWG08, HTWG11, Int03, Int05a, Int05b, Mar05, MR04, WLR05, Wil00, Ano03a, Bro09, HWG04, HWG06, Jon04, KR18, Ses08, Swe06, TCH10, Vic04, Wag02]. languages [Ano04b, Cro01, GS05, Hun03, Int06, MSK09]. Large [CDP16, CHP^{+}08]. large-scale [CHP^{+}08]. last [YH10]. Lead [RIS^{+}17]. Lead-Acid [RIS^{+}17]. Learning [Lib02a, LM06, LM08, MSK09]. Legacy [BHP^{+}01]. Lehrgang [Mös06]. lernen [FH06]. lessons [Kic04]. lets [Ano04a]. Letters [BHP^{+}01]. Level [Bui02, BFGS05, LRT^{+}18]. Libraries [BHP^{+}01, Int05b]. Library [AHG^{+}04, Bal06, GDT^{+}02, Fax08, AHG^{+}04]. lifetime [HBM^{+}06]. Lightweight [SD12]. limit [Ano04b]. linearizability [LTZ14]. Lines [PS17]. linking [DLE06]. LINQ [Nau08a, AA08b, Hum06, Mag10]. Lisp [Kic04]. Lite [KSD12]. Little [Kic04, Wag02]. Loader [BHP^{+}01]. lock [KPS09]. lock-free [KPS09]. Long [Kic04]. Look [Ano01, Hun03, Ste02b, HSF03, HSH04]. Lost [BMT07]. lot [Cro01, Hun03].

Machines [G^{+}01, JK13]. macros [Kic04]. Mainsoft [Ano04a]. mainstream [Swe06]. majors [GD07]. Making [McC04b]. managed [CBGM12, NBV14]. management [KCP^{+}17]. Manual [GDT^{+}02, KCP^{+}17]. Many [GBJ^{+}13, Cro01, GBJ^{+}10, GBJ^{+}12, Kic04]. Many-Body [GBJ^{+}13, GBJ^{+}10, GBJ^{+}12]. Mapping [Nau08a]. Margus [FV09]. Marketing [Lut03]. markets [Wor04]. Matlab [MP18]. MCAD [Ano02, Ano03b]. MCAD/MCSD [Ano02, Ano03b]. MCSD [Ano02, Ano03b]. mechanism [KPS09]. Mehta [Nau08a]. Members [McC03d]. memory [AHM09, KCP^{+}17, ZHU^{+}10]. Merlin [HBM^{+}06]. metaprogramming [Kic04]. metrics [SSK13]. Metro [Gri12b]. Microbenchmarking [Bru05]. Microsoft [Pel06b, Ano02, Ano03a, Ano03b, Ano04a, Ano04b, Bak06, BD05, Bal06, Bra06, Fox04, Hof06, Hol04, Hun03, Mar05, Pel05, Pel06a, PSN^{+}02, Pet02, RHSvS02, SJNH02, Sha06b, Sha06a, Wat10b]. Migration [Gro22]. Minimizing [TWBS09]. model [DLE06, JPS^{+}08, RRY10, Stä05, FV09].

Model-based [FV09]. models [BRE06b]. Modern [BCF04, JK13]. modular [BFGS05]. monad [SWS^{+}13]. Monitoring [RIS^{+}17]. monitors [KPPÉR06]. Mono [DB04]. Moronic [Lut03]. MorphJ [HS08]. move [Ano04a]. MPI [WLR05]. MS [RHSvS02]. MSIL [LN04]. Multicast [MA03]. MultiJava [CMLC06]. multilevel [JK13]. Multiple [RIS^{+}17]. Murach [ML04, Mur06]. Mutating [PVR^{+}20]. Mutation [DS12, DR12]. mutators [MSLL07].

neuromorphic [HNTL12]. News [Bar01]. Next [ACM18, Swe06]. No [Ano17]. notebook [DB04, Lib05b]. novice [CAGH04, HRA^{+}05, Mac05a]. null [BNK^{+}07]. number [Ano04b]. Numerics [AHG^{+}04]. Nutshell [Alb22, AAD07, AA10a, DAN02a, DAN03].

Object [DS12, DR12, Nau08a, RHSvS02, Sax03, BRE06b, CHP^{+}08, FX07, HBM^{+}06, JPS^{+}08, KR18, MSK09, SA09, TCH10]. Object-Oriented [DS12, DR12, RHSvS02, Sax03, BRE06b, CHP^{+}08, JPS^{+}08, KR18, TCH10]. Objects [Lho06, BP04, Fos05, JJOA07, Mag10]. objektorientiert [BRE06b]. objektorientierten [BRE06b]. off [AHM09]. off-the-shelf [AHM09]. offering [Kic04]. offers [Ano04a, Ano05a]. often [Hun03]. OOP [Nod04, RHSvS02]. OpenGJK [MP18]. OpenML [Bar01]. operating [YH10]. Operators [DR12, McC04a]. optimization [MGM^{+}06]. optimizing [CHP^{+}08]. Oracle [Lut03]. order [SWS^{+}13]. Oriented [DS12, DR12, BRE06b, CHP^{+}08, JPS^{+}08, KR18, MSK09, RHSvS02, SA09, Sax03, SSRW02, TCH10]. origin [BNK^{+}07]. orthodox [JJOA07]. Outhouse [Ste02a]. Overloaded [McC04a]. own [Ruv04].

paced [Ano02, Ano03b]. package [AK09]. Palo [ACM01]. Panel [G^{+}01]. paperback [FV09]. Paradigm [Ste02a]. Paradigms [Swa01]. Parallel [Mag10]. parallelism [GPP^{+}12, PRV10]. Parameterized [CKL^{+}13]. parameters [Tit13]. parsing [Sax03]. Partial [PS17, PGBG12]. passing [Fos05]. path [AZLY18]. path-based [AZLY18]. Patients [CDP16]. Patterns [Lut03, Bis07, Coo03, Met04, Nil06, Rus08, Lut03]. people [SSRW02]. people-oriented [SSRW02]. Performance [DS12, SA09, CBGM12, G^{+}01]. Perl [SKS08, Cro01]. Perry [Ano17]. perspective [PW01, Swe06]. PHP [SKS08]. Platform [Gri02, LN04, McS10, Shu05, Tro01, Tro03, Tro05, Tro10]. platforms [Ano04a]. PLINQ [Mag10]. Pocket [DAN02b, Moj02, AA08a, AA08b, AA10b]. point [Mös04]. policies [KPPÉR06]. policy [JK13]. Population [CDP16]. Posit [LRT^{+}18]. power [CBGM12]. pp. [FV09]. Practical [BD05, HP09, MA03, WT10, dCP05]. practices [BD05, VB02]. Pragmatic [Wil03]. precisely [SH04]. predicting [AZLY18]. Prediction [CDP16, ZWZ^{+}14]. Presentation [Mac08, Nau08b]. preserving [CHP^{+}08, McS10]. Press [Ano17]. prevalence [Ano03a]. principled [Kic04]. principles [BH04b]. Privacy [McS10]. privacy-preserving [McS10]. Pro [Nau08a, Nau08b, Tro05, Mac05b, Mac08, Tro10]. Pro.NET [Mac06]. probabilistic [ZWZ^{+}14]. Problem [Doy10, JM06, BH02, Doy08, Hof03, JF10]. Problem-Solution [JM06, JF10]. procedural [BRE06a]. Proceedings [ACM01]. processing [NBV14]. Product [PS17, Ano04a]. Professional [ERH10, NEG^{+}06, NEG^{+}10, Rob02, RAC^{+}03, RNG^{+}04, CAGH04, HRA^{+}05, Mac05a, Mac05b]. professionals [R^{+}02]. Profiles [Int05b]. Program [Doy10, McC03b, AZLY18, Ano05b, Dei02, DD06b, DD11, Doy08, KSD12, Lew06, MGM^{+}06, Pel05]. Programm [BRE06b]. programmable [PRV10]. Programmeercursus [RHSvS02, SJNH02]. Programmer [Gri02, SG01, Wil01, D^{+}03, GWH05, Mac04]. Programmeren [AWvS^{+}02, GN01, PSN^{+}02]. Programmering [Arc02]. programmers [DD06a, DD09, GN01, MA03, VB02, dCP05]. programmeurs [GN01]. programmieren [BRE06b, FH06, Pel06b, BRE06a]. Programmierung [BRE06a]. Programming [ADM02, AMSBA15, Bal06, Bee06, Bui02, Doy10, Eri01, FM02, FJ03, GK03, GAL10, Gri12a, Gri12b, HTWG08, HTWG11, Int06, Ler10, Lib01, LH02, Lib02b, LH03, Lib03, LH04, Lib05a, LM06, LX08, Löw10, Mar05, McK16, Pet02, Ses08, Ste02a, Ste02b, Swa01, WAH03, Wil00, Ano04b, Bak06, BRE06a, BRE06b, BH02, Bey01, BH04b, Bra06, Buo03, Buo05, CAGH04, Cha04, Doy08, DLW05, H^{+}00, HWG04, HWG06, Hof03, JPS^{+}08, Kim02, MSK09, Pen05, PSN^{+}02, RRY10, RHSvS02, Sel04, SG04, SJNH02, Swe06, Tho02, Vic04, YD04, GN01]. Programs [DR12, Lut03, BRE06b, JPS^{+}08, PVR^{+}20, SK17, Sub05, SWS^{+}13, TWBS09, Wat10a, ZHU^{+}10]. Project [Wan11]. proneness [GC15]. Properties [McC03d, AZLY18]. proposed [BMT07]. protection [AHM09]. provide [Kic04]. proving [Fru10]. Proxies [Bar03]. prozedural [BRE06a]. public [H^{+}00]. Publications [Bee06]. published [FV09]. Python [SKS08].

Quality [DR12]. queries [McS10, MP18]. query [NBV14]. queuing [KPPÉR06]. quick [DAN02a, DAN03].

rationale [CMLC06]. reactions [GBJ^{+}12, GBJ^{+}13]. Ready [LL03a, LL03b]. Real [AMSBA15, LL03a, LL03b, RIS^{+}17, Wat10a, KPHV11, PPS08, PZM^{+}10]. Real-Time [LL03b, RIS^{+}17, KPHV11, PPS08, PZM^{+}10]. Recipes [JM06, JF10]. Record [BHP^{+}01]. recursive [Sax03]. refactoring [GC15]. Reference [AHG^{+}04, Alb22, DAN02b, GDT^{+}02, Moj02, AA08a, AA08b, AA10b, DAN02a, DAN03, GPP^{+}12, Her02, Sch06a, Sch10]. Refinement [JLP^{+}14, KPPÉR06]. Reflection [McC04d, HS08]. Reflexes [SPP^{+}10]. Reguläre [SKS08]. Relational [Nau08a]. relativistic [GBJ^{+}10, GBJ^{+}12, GBJ^{+}13]. Released [Bar01]. Reliable [MP18]. Relief [Bar01]. Repairs [Ste02b]. Replace [Reg02a]. reporting [BNK^{+}07]. representation [AZLY18, MGM^{+}06]. research [HP04]. Resorts [ACM18]. resources [Ano04b]. responsive [SPP^{+}10]. ResumableFunctions [Lau17]. Reuse [BS04, AK09]. Review [Ano17, Cir01, Nau08a, Nau08b, FV09]. Richtig [FH06]. Ruby [SKS08]. runtime [AK09, BS03]. runtimes [NBV14]. RUST [CDP16].

Safe [PRV10, YH10, Fek08, GPP^{+}12, HS08, KCP^{+}17]. Safer [FLF08]. safety [Fru10]. Sams [Jon02, Jon04, Fox04, Fox06, Fox08]. Says [Bar01]. scale [CHP^{+}08]. scanning [KPS09]. Scheduling [KPHV11]. Schema [Lut03]. Schemas [Lut03]. SCHISM [PZM^{+}10]. Schulte [FV09]. Science [Bar01, GD07]. Scientific [GDT^{+}02, GK03]. Scratch [AMSBA15]. Search [Lut03]. secrets [Bra06]. Security [Ano05b, FJ03, MJ03]. selection [SVY09]. Self [Ken05, Ano02, Ano03b]. self-paced [Ano02, Ano03b]. Self-Teaching [Ken05]. semantics [BFGS05]. sequence [ZWZ^{+}14]. Serialization [McC05]. Server [Bui02, Rei02, Ano03b]. Services [FM02, Ano03a, Ano03b, Bal03, BW02, Löw10]. set [ABE^{+}02]. Setting [BHP^{+}01]. Shared [BHP^{+}01]. sharp [Hun03, Lau17]. shelf [AHM09]. SIGCSE [Reg02b]. Simple [KCP^{+}17]. Simulation [CDP16]. Simultaneous [LRT^{+}18]. Singapore [ACM18]. single [JK13]. smaller [TWBS09]. SML [GS05]. Socket [Bui02]. Socket-Level [Bui02]. sockets [MA03]. Software [LRT^{+}18, PS17, Wan11, CBGM12, FV09, GC15, Lew06, SA09, Ses08, SSRW02, LL03b]. Softwareentwicklung [Mös06]. Solution [JM06, BH02, Hof03, JF10]. Solutions [Wat10a, Lew06, MP18]. Source [Ano05b, BHP^{+}01, SSK13]. Source-Code [BHP^{+}01]. space [MP18]. Spec [Bar01]. Spec# [BFL^{+}11]. specific [Wag05, Wag10]. Specification [BFL^{+}11, BS04, C#02, ECM01, ECM02a, ECM05, ECM06, ECM02b, Int03, Stä05]. speculative [PRV10]. SQL [Ano05b]. SQLite [New05]. SSA [MGM^{+}06]. stack [KPS09]. Standard [AHG^{+}04, C#02, DR12, MR04, JPS07]. Star [Lut03, Lut03]. Starch [Ano17]. start [Ano03a, Hol04]. Started [LM06]. Starter [Bar06, Fox06, Fox08]. Starting [CG05]. Static [McC03d, SK17, XJC09, HS08, SD12]. stealing [Fax08]. Step [Sha06a, Sha06b]. Straight [BHP^{+}01]. Stream [KBPS17]. streams [CS06]. strengths [Ano04b]. stress [WNG03]. strong [AHM09]. structure [SB07]. structured [BRE06a]. strukturiert [BRE06a]. students [Fek08]. Studio [CAGH02]. Study [BS04, KPPÉR06, PPS08]. Stupidity [Lut03]. style [BGM06, Lau17]. Suite [Ano05b]. Sun [Ano04b]. support [Ano04b]. supporting [AK09]. syntax [SSK13]. System [DS12, RIS^{+}17, YH10]. System.XML [HSF03]. Systems [LN04, LL03b, GBJ^{+}10, GBJ^{+}12, GBJ^{+}13, RRY10, SD12, SA09].

Talent [Bar01]. tasks [SPP^{+}10]. teach [Fox04, Fox06, Fox08, Jon02, Jon04]. Teaching [Fek08, Ken05]. Tech [Lut03]. technique [SSK13]. technology [Int03, Int05a, Int05b, Int06, KSD12, Kic04]. templates [AK09]. Testing [DS12, FV09]. Text [Lut03, Mey02]. their [MSLL07]. things [McK16, RIS^{+}17]. thread [Fek08, Stä05]. thread-safe [Fek08]. threading [TFRS03]. three [MP18]. three-dimensional [MP18]. throughout [Gou08]. tight [Ano04b]. Time [LL03a, LL03b, RIS^{+}17, Wil01, KPHV11, PPS08, PZM^{+}10, STW15]. together [STW15]. tolerance [RV13]. tolerant [PZM^{+}10]. Tool [CDP16, SSK13]. tower [Reg02b]. TR [Int05b]. traces [HBM^{+}06]. Tracking [BNK^{+}07]. Traditional [GS05]. training [Ano02, Ano03b]. Transactional [AHM09, ZHU^{+}10]. translation [BMT07]. Translator [LN04]. Transposing [KR18]. tree [SSK13]. Trio [SG01]. try [Ano04b]. tuples [Tit13]. tutorial [Coo03]. Two [Ano01]. Type [CHP^{+}08, PGBG12, ABE^{+}02, Fru10, Tit13, YH10, LRT^{+}18]. Type-directed [PGBG12]. Type-preserving [CHP^{+}08]. type-safe [YH10]. typed [TCH10]. Types [McC03a, FX07, KR18].

uncertain [McK16]. undefined [BNK^{+}07]. uniprocessors [KPHV11]. Uniqueness [GPP^{+}12]. unleashed [Hof06, May08]. Unlocking [Bra06]. unsafe [FLF08]. Unum [LRT^{+}18]. Use [McC04b, PS17, Ano05a, ZHU^{+}10]. Using [LRT^{+}18, Lut03, McC03d, McC04c, McC04d, Wan11, WLR05, AHM09, BH04b, BW02, KMM^{+}02, Mag10, Ruv04, SSK13, SSRW02]. uur [FHCM^{+}02].

v [HSH04]. v.2.0 [HSF03]. v03 [GBJ^{+}13]. v2.0 [Bra06]. value [BNK^{+}07]. VB [ERH10, GS05]. VB.NET [Moj02, RHSvS02, Ruv04, Sub05]. Veanes [FV09]. verifiable [MGM^{+}06]. verification [BFL^{+}11, HP09, KSD12, Stä05, YH10]. Verified [JLP^{+}14]. Verifying [SWS^{+}13]. version [GBJ^{+}13]. Versus [Ead01]. via [Nod04, RV13, Ric06, Ric10]. Views [Bar01, BH04b]. Vijay [Nau08a]. violations [LTZ14]. Virgil [Tit13]. Virtual [G^{+}01]. Visual [Ano02, Ano03b, Bak06, BD05, Bal06, FH06, Fox04, Hof06, Hol04, Mar05, Pel05, Pel06a, Pel06b, RHSvS02, SK08, SJNH02, Sha06b, Sha06a, Wat10b, Ano04a, BD05, Bar06, CAGH02, CAGH04, DD06b, DD11, Fox06, Fox08, Gro22, JM06, JF10, Ken05, Lib05b, Oni03b, PH03, Rus08, Sch06b, SG04, Tel06, UKH^{+}04, Vic04, WEG^{+}02, W^{+}05, W^{+}06]. Volume [AHG^{+}04, BS03]. vom [BRE06b]. voor [GN01, R^{+}02].

Warehousing [Lut03]. Warps [Wil01]. Wave [Ste02b]. Way [Kic04]. ways [Sub05, Wag05, Wag10]. WCF [Löw10]. Weaving [CSH03]. Web [Ano03a, Ano03b, Bak02, Bal03, BW02, CAGH02, CAGH04, FM02, Gri12b, Lut03]. Website [BH02, Ruv04]. which [Ano04b]. Windows [Ano02, Nau08b, Gri12b, LH04, Mac08, Mac06, PSN^{+}02, Pet02, Sel04, SG04]. Windows-based [Ano02]. Winners [Bar01]. within [TWBS09]. without [WNG03]. Wolfram [FV09]. Wool [Fax08]. work [Fax08]. workbench [MSK09]. Working [McC03e, McC04e, McC05]. World [ACM18, McC03b, McK16, Wil03]. WPF [Nau08b, Mac08, Mac10].

XML [Ano03b, FL02].

yang [CBGM12]. Years [Lut03, Kic04]. yin [CBGM12]. You’d [ABE^{+}02]. yourself [Fox04, Fox06, Fox08, Jon02, Jon04].

zonder [WNG03]. zum [BRE06b]. Zuse [BHP^{+}01].

References

Albahari:2008:CPR

[AA08a] Joseph Albahari and Ben Albahari. C# 3.0 pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2008. ISBN 0-596-51922-2 (paperback). xi + 230 pp. LCCN QA76.73.C154 A4185 2008.

Albahari:2008:LPR

[AA08b] Joseph Albahari and Ben Albahari. LINQ pocket reference. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51924-9 (paperback). ix + 161 pp. LCCN QA76.73.L228 A43 2008. URL http://www.loc.gov/

catdir/toc/fy0805/2008299898.html.

Albahari:2010:CN

[AA10a] Joseph Albahari and Ben Albahari. C# 4.0 in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2010. ISBN 0-596-80095-9. xviii + 1025 pp. LCCN ????

Albahari:2010:CPR

[AA10b] Joseph Albahari and Ben Albahari. C# 4.0 pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2010. ISBN 1-4493-9401-9 (paperback). ???? pp. LCCN ????

Albahari:2007:CN

[AAD07] Joseph Albahari, Ben Albahari, and Peter Drayton. C# 3.0 in a nutshell. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2007. ISBN 0-596-52757-8 (paperback). xviii + 838 pp. LCCN QA76.73.C154 D73 2007; QA76.73.C154; QA76.73.C154 D73

2007eb. URL http://www.oreilly.com/catalog/9780596527570.

Alm:2002:YBS

[ABE^{+}02] Jacob Alm, Robert Baber, Shauna Eggers, Christopher O’Toole, and Abin Shahab. You’d better set down for this!: creating a set type for CS1 & CS2 in C#. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):14–18, September 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

ACM:2001:PAJ

[ACM01] ACM, editor. Proceedings of the ACM 2001 Java Grande/ISCOPE Conference: Palo Alto, Calif., June 2–4, 2001. ACM Press, New York, NY 10036, USA, 2001. ISBN 1-58113-359-6. LCCN QA76.9.O35 A26 2001.

ACM:2018:CNG

[ACM18] ACM, editor. Conference for Next Generation Arithmetic (Resorts World Convention Centre, Singapore): (CoNGA ’18). ACM Press, New York, NY 10036, USA, 2018.

Albahari:2001:CE

[ADM01] Ben Albahari, Peter Drayton, and Brad Merrill. C# Essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, February 2001. ISBN 0-596-00079-0. xiv + 199 pp. LCCN QA76.73.C154 A418 2001.

US$19.95. URL http://www.oreilly.com/catalog/csharpess.

Albahari:2002:CEP

[ADM02] Ben Albahari, Peter Drayton, and Brad Merrill. C# Essentials: Programming the .NET Framework. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00315-3. xii + 202 pp. LCCN QA76.73.C154 A418 2002. US$24.95.

URL http://safari.oreilly.com/0596003153;   http://www.
oreilly.com/catalog/csharpess2.

Abrams:2004:NFS

[AHG^{+}04] Brad Abrams, Anders Hejlsberg, Brian Grunkemeyer, Joel Marcey, Kit George, Krzysztof Cwalina, and Jeffrey Richter. .NET Framework Standard Library Annotated Reference. Volume 1: Base Class Library and Extended Numerics Library. Microsoft .NET development series. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-321-15489-4 (hardcover). xxvi + 528 pp. LCCN

QA76.76.M52 A27 2004. URL http://www.aw-bc.com/catalog/
academic/product/0,1144,0321154894,00.html; http://www.
loc.gov/catdir/toc/ecip0411/2003024327.html. Foreword by

Joel Marcey.

Abadi:2009:TMS

[AHM09] Martín Abadi, Tim Harris, and Mojtaba Mehrara. Transactional memory with strong atomicity using off-the-shelf memory protection hardware. ACM SIGPLAN Notices, 44(4):185–196, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Axelsen:2009:GPT

[AK09] Eyvind W. Axelsen and Stein Krogdahl. Groovy package templates: supporting reuse and runtime adaption of class hierarchies. ACM SIGPLAN Notices, 44(12):15–26, December 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Albahari:2022:CND

[Alb22] Joseph Albahari. C# 10 in a Nutshell: the Definitive Reference. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2022. ISBN 1-0981-2195-3 (paperback), 1-0981-2191-0 (e-book), 1-0981-2192-9 (e-book). xv + 1041 pp. LCCN QA76.73.C154 A5325 2022.

Armoni:2015:SRP

[AMSBA15] Michal Armoni, Orni Meerbaum-Salant, and Mordechai Ben-Ari. From scratch to “real” programming. ACM Transactions on Computing Education, 14(4):25:1–25:??, February 2015. CODEN ???? ISSN 1946-6226.

Anonymous:2001:LTC

[Ano01] Anonymous. A look at two C# books. C/C++ Users Journal, 19 (12):??, December 2001. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2002:MMS

[Ano02] Anonymous. MCAD/MCSD self-paced training kit: developing Windows-based applications with Microsoft Visual Basic .NET and Microsoft Visual C# .NET, Exams 70-306 and 70-316. Microsoft Press, Redmond, WA, USA, 2002. ISBN 0-7356-1533-0. xxvii + 541 pp. LCCN CD-ROM; 005; QA76.3 .M32556 2002; DVD.

Anonymous:2003:JHS

[Ano03a] Anonymous. Java’s head start adoption of Microsoft’s C# language for building Web services is hindered by the prevalence of Java. Information Week, 966:57, 2003. CODEN INFWE4. ISSN 8750-6874.

Anonymous:2003:MMS

[Ano03b] Anonymous. MCAD/MCSD self-paced training kit: Developing XML Web services and server components with Microsoft Visual Basic .NET and Microsoft Visual C# .NET: Exams 70-310 and 70-320. Microsoft Press, Redmond, WA, USA, 2003. ISBN 0-7356-1586-1. xxix + 651 pp. LCCN Pucbb; CD-ROM; 005. UK£45.99.

Anonymous:2004:BBM

[Ano04a] Anonymous. Building bridges: Mainsoft offers a product that lets companies move applications built in Microsoft’s Visual Basic and C# to Java platforms. Information Week, 977:30–31, 2004. CODEN INFWE4. ISSN 8750-6874.

Anonymous:2004:CGH

[Ano04b] Anonymous. C# goes head to head with Java: When resources are tight, companies try to limit the number of programming languages they support. Microsoft’s C# and Sun’s Java both have strengths, so which do you choose? Computer Weekly, pages 48–50, November 9, 2004. CODEN ???? ISSN 0010-4787.

Anonymous:2005:COE

[Ano05a] Anonymous. C# offers an easy-to-use alternative to Java. Computer Weekly, page 30, January 18, 2005. CODEN ???? ISSN 0010-4787.

Anonymous:2005:SAS

[Ano05b] Anonymous. Security audits for source code fortify source code: Analysis Suite 3.0 brings deep, intelligent security analysis to C, C++, C#, Java, JSP, and SQL program code. InfoWorld, 27(3): 24–25, 2005. CODEN INWODU. ISSN 0199-6649.

Anonymous:2017:BRG

[Ano17] Anonymous. Book review: Gray Hat C#, by Brandon Perry. No Starch Press. ISBN 978-1-59327-759-8. Network Security, 2017(7): 4, July 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-

9371 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1353485817300685.

Archer:2002:PCT

[Arc02] Tom Archer. Programmering i C#. Gyldendal akademisk, Oslo, Norway, 2002. ISBN 82-05-30366-5. xx + 378 pp. LCCN ????

Archer:2002:IC

[AW02] Tom Archer and Andrew Whitechapel. Inside C#. Microsoft Press, Redmond, WA, USA, second edition, 2002. ISBN 0-7356-1648-5 (paperback). xxxiii + 873 pp. LCCN QA76.73.C154 A73 2002.

Archer:2002:PMC

[AWvS^{+}02] Tom Archer, Andrew Whitechapel, Bas van Spijker, Wout Natter, et al. Programmeren met C#. Academic Service, Schoonhoven, The Netherlands, second edition, 2002. ISBN 90-395-1927-7. xxxi + 874 pp. LCCN ????

Alon:2018:GPB

[AZLY18] Uri Alon, Meital Zilberstein, Omer Levy, and Eran Yahav. A general path-based representation for predicting program properties. ACM SIGPLAN Notices, 53(4):404–419, April 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bakharia:2002:CFE

[Bak02] Aneesha Bakharia. C#: fast and easy Web development. Premier, Rocklin, CA, USA, 2002. ISBN 1-931841-05-5 (paperback). xx + 553 pp. LCCN QA76.73.C154 B35 2002eb.

Bakharia:2006:MVC

[Bak06] Aneesha Bakharia. Microsoft Visual C# 2005 Express Edition: programming for the absolute beginner. For the absolute beginner. Thomson Course Technology PTR, Boston, MA, USA, 2006. ISBN 1-59200-818-6 (paperback). xix + 259 pp. LCCN QA76.73.C154 B35

2006. URL http://link.library.utoronto.ca/eir/EIRdetail.
cfm?Resources\_\_ID=370659&T=; resource.

Ballinger:2003:NWS

[Bal03] Keith Ballinger. .NET Web services: architecture and implementation. Microsoft .NET development series. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-321-11359-4. xix + 323 pp. LCCN TK5105.888 .B355 2003.

Balena:2006:PMV

[Bal06] Francesco Balena. Programming Microsoft Visual C# 2005: The Base Class Library. Microsoft Press, Redmond, WA, USA, 2006. ISBN 0-7356-2308-2. ???? pp. LCCN ????

Baran:2001:NVA

[Bar01] Nicholas Baran. News and views: Anonymity and the Internet; is industry hogging computer science talent?; relief from acronyms; OpenML spec released; C# not just a Java killer, says anders; and the winners are . . .. Dr. Dobb’s Journal of Software Tools, 26(7):18, July 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/.

Barrett:2003:DPJ

[Bar03] Tom Barrett. Dynamic proxies in Java and .NET. Dr. Dobb’s Journal of Software Tools, 28(7):18, 20, 22, 24, 26, July 2003. CODEN

DDJOEB. ISSN 1044-789X. URL http://www.ddj.com/ftp/2003/
2003_07/dyproxy.zip.

Barker:2006:VCE

[Bar06] F. Scott Barker. Visual C# 2005 Express Edition Starter Kit. Wrox Press, Chicago, IL, USA, 2006. ISBN 0-7645-8955-5. 336 (est.) pp. LCCN ????

Bayer:2007:CCH

[Bay07] Jürgen Bayer, editor. Das C#-2005-Codebook. (German) [The C# Codebook]. eBook. Addison-Wesley, München, Germany, 2007. ISBN 3-8273-6118-4. ???? pp. LCCN ????

Bayer:2008:CCH

[Bay08] Jürgen Bayer. Das C# 2008 Codebook. (German) [The C# Codebook]. Addison Wesley, München, Germany, 2008. ISBN ???? ???? pp. LCCN ????

Benton:2004:MCA

[BCF04] Nick Benton, Luca Cardelli, and Cédric Fournet. Modern concurrency abstractions for C#. ACM Transactions on Programming Languages and Systems, 26(5):769–804, September 2004. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Balena:2005:PGB

[BD05] Francesco Balena and Giuseppe Dimauro. Practical guidelines and best practices for Microsoft Visual Basic and Visual C# developers. Microsoft Press, Redmond, WA, USA, 2005. ISBN 0-7356-2172-1. xxxiv + 570 pp. LCCN QA76.73.B3 B344 2005.

Beebe:2006:BPAa

[Bee06] Nelson H. F. Beebe. A bibliography of publications about the C# programming language. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA,

March 15, 2006. 19 pp. URL https://www.math.utah.edu/pub/
tex/bib/index-table-c.html#csharp.

Beyer:2001:CCP

[Bey01] Derek Beyer. C# COM+ programming. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 2001. ISBN 0-7645-4835-2 (paperback). 408 pp. LCCN QA76.73.C154 B49 2001b.

Borger:2005:HLM

[BFGS05] Egon Börger, Nicu G. Fruja, Vincenzo Gervasi, and Robert F. Stärk. A high-level modular definition of the semantics of C#. Theoretical Computer Science, 336(2–3):235–284, May 26, 2005. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.only4gurus.net/miscellaneous/cs03.pdf.

Barnett:2011:SVS

[BFL^{+}11] Mike Barnett, Manuel Fähndrich, K. Rustan M. Leino, Peter Müller, Wolfram Schulte, and Herman Venter. Specification and verification: the Spec# experience. Communications of the Association for Computing Machinery, 54(6):81–91, June 2011. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Baldwin:2006:ECS

[BGM06] Kenneth Baldwin, Andrew Gray, and Trevor Misfeldt. The elements of C# style. Cambridge University Press, Cambridge, UK, 2006. ISBN 0-521-67159-0 (paperback). x + 146 pp. LCCN

QA76.73.C154 B36 2006. URL http://www.loc.gov/catdir/
enhancements/fy0633/2006008307-d.html;   http://www.loc.

gov/catdir/enhancements/fy0642/2006008307-t.html;   http:
//www.loc.gov/catdir/enhancements/fy0732/2006008307-b.html.

Bellinaso:2002:ANW

[BH02] Marco Bellinaso and Kevin Hoffman. ASP.NET Website programming: problem, design, solution. Programmer to programmer. Wrox Press, Chicago, IL, USA, C# edition, 2002. ISBN 0-7645-4377-6 (paperback). ix + 540 pp. LCCN TK5105.8885.A26 B45 2002b.

Bishop:2004:CC

[BH04a] J. M. (Judith Mary) Bishop and Nigel Horspool. C# concisely. Pearson/Addison Wesley, Harlow, UK, 2004. ISBN 0-321-15418-5 (paperback). xxiii + 442 pp. LCCN QA76.73.C154 B57 2004. US$29.99.

Bishop:2004:DPG

[BH04b] Judith Bishop and Nigel Horspool. Developing principles of GUI programming using views. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):373–377, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bonzini:2001:LHG

[BHP^{+}01] Paolo Bonzini, Stuart Halloway, John Penry, Oluseyi Sonaiya, Bruce E. Hogman, Greg Bissell, Michael Hobbs, and Ben Laurie. Letters: Huge GCC executables; Java class loader; Department of Dumb Ideas; setting the record straight; the legacy of C#; DHTML source-code correction; shared libraries aren’t all bad; Zuse and Intel. Dr. Dobb’s Journal of Software Tools, 26(8):10, 12, August 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.ddj.com/.

Bishop:2005:EIJ

[BHW05] Judith Bishop, R. Nigel Horspool, and Basil Worrall. Experience in integrating Java with C# and .NET. Concurrency and Computation: Practice and Experience, 17(5–6):663–680, April/May 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bishop:2004:CGC

[Bis04] David Bishop. A complete guide to C#. Jones and Bartlett, Boston, MA, USA, 2004. ISBN 0-7637-2249-9. xvii + 624 pp. LCCN QA76.73.C154 B55 2004.

Bishop:2007:CDP

[Bis07] J. M. (Judith Mary) Bishop. C# design patterns. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52773-X (paperback). xxi +

290 pp. LCCN Internet. URL http://www.oreilly.com/catalog/
9780596527730.

Bierman:2007:LTF

[BMT07] Gavin M. Bierman, Erik Meijer, and Mads Torgersen. Lost in translation: formalizing proposed extensions to C#. ACM SIGPLAN Notices, 42(10):479–498, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bond:2007:TBA

[BNK^{+}07] Michael D. Bond, Nicholas Nethercote, Stephen W. Kent, Samuel Z. Guyer, and Kathryn S. McKinley. Tracking bad apples: reporting the origin of null and undefined value errors. ACM SIGPLAN Notices, 42(10):405–422, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bo:2002:C

[Bo02] Brinch Bo. C#. Udvikling af .NET-applikationer med C#. IDG, Valby, Denmark, 2002. ISBN 87-7843-519-6. 212 pp. LCCN ???? DKR 249,00.

Barker:2004:BCO

[BP04] Jacquie Barker and Grant Palmer. Beginning C# objects: from concepts to code. Apress, Berkeley, CA, USA, 2004. ISBN 1-59059-360-X. xxvii + 819 pp. LCCN QA76.73.C154 B375 2004b. URL

http://www.columbia.edu/cgi-bin/cul/resolve?clio4997313.

Braden:2006:UMC

[Bra06] Richard P. Braden. Unlocking Microsoft C# v2.0 programming secrets. Wordware Publishing, Plano, TX, USA, 2006. ISBN 1-55622-097-9 (paperback). xx + 375 pp. LCCN QA76.65 B73 2006. URL

http://www.loc.gov/catdir/toc/ecip055/2004030069.html.

Balzert:2006:CEP

[BRE06a] Helmut Balzert, Michael Ricken, and Thomas Erler. CSharp 2 — der Einstieg in die Programmierung: strukturiert und prozedural programmieren. (German). [CSharp 2 — the entrance into programming: structured and procedural programming. W3L-Verlag, Herdecke, Germany, 2006. ISBN 3-937137-18-1. LCCN ????

Balzert:2006:COP

[BRE06b] Helmut Balzert, Michael Ricken, and Thomas Erler. CSharp 2 — objektorientiert programmieren: vom objektorientierten Analysemodell bis zum objektorientierten Programm. (German) [CSharp 2 – object-oriented programming: from object-oriented analysis models to object-oriented programs. W3L-Verlag, Herdecke, Germany, 2006. ISBN 3-937137-19-X. LCCN ????

Brosgol:2009:ICL

[Bro09] Ben Brosgol. An introduction to the C# language and .NET infrastructure. ACM SIGADA Ada Letters, 29(3):3–4, December 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bruckschlegel:2005:MCC

[Bru05] T. Bruckschlegel. Microbenchmarking C++, C#, and Java. C/C++ Users Journal, 23(7):14–21, 2005. CODEN CCUJEX. ISSN 1075-2838.

Box:2003:ENC

[BS03] Don Box and Chris Sells. Essential .NET: Volume 1: The Common Language Runtime. Microsoft .NET development series. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-73411-7 (paperback). xxiii + 405 pp. LCCN QA76.76.M52 B69 2003. Foreword by James S. Miller.

Borger:2004:EAS

[BS04] E. Borger and R. F. Stark. Exploiting abstraction for specification reuse. the Java/C# case study. Lecture Notes in Computer Science, 3188:42–76, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buis:2002:SLS

[Bui02] Paul Buis. Socket-level server programming and .NET. Dr. Dobb’s Journal of Software Tools, 27(9):25–26, 28, 30, 32, September 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/ftp/2002/2002_09/tcpecho.txt;   http://www.
ddj.com/ftp/2002/2002_09/tcpecho.zip.

Buono:2003:CGP

[Buo03] Salvatore A. Buono. C# and game programming: a beginner’s guide. A. K. Peters, Ltd., Wellesley, MA, USA, 2003. ISBN 1-56881-193-4. xxiii + 556 pp. LCCN QA76.73.C154 B85 2003.

Buono:2005:CGP

[Buo05] Salvatore A. Buono. C# and game programming: a beginner’s guide. A. K. Peters, Ltd., Wellesley, MA, USA, second edition, 2005. ISBN 1-56881-236-1 (paperback). xxiii + 505 pp. LCCN QA76.73.C154 B85 2005.

Bustos:2002:BNW

[BW02] Joe Bustos and Karli Watson. Beginning .NET Web services using C#. Programmer to programmer. Wrox Press, Chicago, IL, USA, 2002. ISBN 1-86100-757-4. v + 342 pp. LCCN ????

Csharp:2002:SEC

[C#02] ECMA International, 114 Rue du Rhône, CH-1204 Geneva, Switzerland. Standard ECMA-334: C# Language Specification, second edi-

tion, December 2002. URL http://www.ecma-international.org/
publications/files/ECMA-ST/Ecma-334.pdf.

Cazzulino:2002:BCW

[CAGH02] Daniel Cazzulino, Victor Garcia Aprea, James Greewood, and Chris Hart. Beginning C# Web applications: with Visual Studio .NET. Programmer to programmer. Wrox Press, Chicago, IL, USA, 2002. ISBN 1-86100-732-9 (paperback). ix + 598 pp. LCCN ???? US$28.99.

Cazzulino:2004:BVW

[CAGH04] Daniel Cazzulino, Victor Garcia Aprea, James Greenwood, and Chris Hart, editors. Beginning Visual Web programming in C#: from novice to professional. Apress, Berkeley, CA, USA, 2004. ISBN 1-59059-361-8. xxiv + 639 pp. LCCN QA76.625 .B44 2004b. URL

http://www.columbia.edu/cgi-bin/cul/resolve?clio5326019.

Cao:2012:YYP

[CBGM12] Ting Cao, Stephen M. Blackburn, Tiejun Gao, and Kathryn S. McKinley. The yin and yang of power and performance for asymmetric hardware and managed software. ACM SIGARCH Computer Architecture News, 40(3):225–236, June 2012. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ISCA ’12 conference proceedings.

Conway:2003:CCD

[CDH^{+}03] Richard Conway, Teun Duynstee, Ben Hyrman, Roger Rowland, and James Speer, editors. C# class design handbook: coding effective classes. Wrox Press, Chicago, IL, USA, 2003. ISBN 1-59059-257-3, 1-86100-828-7 (paperback). vii + 366 pp. LCCN QA76.73.C154 C743 2003b. US$28.99.

Cimler:2016:CRC

[CDP16] Richard Cimler, Ondřej Doležal, and Pavel Pscheidl. Comparison of RUST and C# as a tool for creation of a large agent-based simulation for population prediction of patients with Alzheimer’s disease in EU. In Computational Collective Intelligence: 8th International Conference, ICCCI 2016, Halkidiki, Greece, September 28–30, 2016. Proceedings, Part II, pages 252–261. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. URL http://

link.springer.com/chapter/10.1007/978-3-319-45246-3_24.

Chegwidden:2005:SC

[CG05] James Chegwidden and Tony Gaddis. Starting out with C#. Scott/James Publishers, El Granada, CA, USA, 2005. ISBN 1-57676-161-4. xvi + 790 pp. LCCN QA76.73.C154 C444 2005. URL ftp:/

/ftp.aw.com/cseng/authors/chegwidden/   ( materials).

Chand:2004:GPG

[Cha04] Mahesh Chand. Graphics programming with GDI+. Addison-Wes-ley, Reading, MA, USA, 2004. ISBN 0-321-16077-0. xliv + 740 pp. LCCN T385 .C4515 2003.

Chen:2008:TPC

[CHP^{+}08] Juan Chen, Chris Hawblitzel, Frances Perry, Mike Emmi, Jeremy Condit, Derrick Coetzee, and Polyvios Pratikaki. Type-preserving compilation for large-scale optimizing object-oriented compilers. ACM SIGPLAN Notices, 43(6):183–192, June 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cirillo:2001:BRC

[Cir01] William Cirillo. Book review: C/C++ Contributing Editors. C/C++ Users Journal, 19(12):??, December 2001. CODEN CCU-JEX. ISSN 1075-2838.

Choi:2013:PAC

[CKL^{+}13] Byoung K. Choi, Donghun Kang, Taesik Lee, Arwa A. Jamjoom, and Maysoon F. Abulkhair. Parameterized activity cycle diagram and its application. ACM Transactions on Modeling and Computer Simulation, 23(4):24:1–24:??, October 2013. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Clifton:2006:MDR

[CMLC06] Curtis Clifton, Todd Millstein, Gary T. Leavens, and Craig Chambers. MultiJava: Design rationale, compiler implementation, and applications. ACM Transactions on Programming Languages and Systems, 28(3):517–575, May 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Cooper:2003:CDP

[Coo03] James W. (James William) Cooper. C# design patterns: a tutorial. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-84453-2 (paperback). 392 pp. LCCN QA76.73.C154 C664 2003. US$34.99.

Crowell:2001:CP

[Cro01] Talbott Crowell. C# and Perl: C# has a lot in common with many languages. Dr. Dobb’s Journal of Software Tools, 26(10):50, 52, 54, October 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/2001/2001_10/csharpperl.txt.

Chilimbi:2006:CCC

[CS06] Trishul M. Chilimbi and Ran Shaham. Cache-conscious coallocation of hot data streams. ACM SIGPLAN Notices, 41(6):252–262, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carlisle:2003:WAN

[CSH03] Martin C. Carlisle, Ricky E. Sward, and Jeffrey W. Humphries. Weaving Ada 95 into the .NET environment. ACM SIGADA Ada Letters, 23(1):22–26, March 2003. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Deitel:2003:CPI

[D^{+}03] Harvey M. Deitel et al., editors. C#: a programmer’s introduction. Deitel developer series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2003. ISBN 0-13-046132-6. xlix + 862 pp. LCCN QA76.73.C154 C14 2003.

Drayton:2002:CND

[DAN02a] Peter Drayton, Ben Albahari, and Ted Neward. C# in a nutshell: a desktop quick reference. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00181-9. xxii + 832 pp. LCCN QA76.73.C154 D73 2002. US$39.95.

URL http://safari.oreilly.com/0596001819;   http://www.
oreilly.com/catalog/csharpnut.

Drayton:2002:CLP

[DAN02b] Peter Drayton, Ben Albahari, and Ted Neward. C# Language Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00429-X. ix + 118 pp. LCCN QA76.73.C155

D73 2003. US$12.95. URL http://www.oreilly.com/catalog/
csharplangpr.

Drayton:2003:CND

[DAN03] Peter Drayton, Ben Albahari, and Ted Neward. C# in a nutshell: a desktop quick reference. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00526-1. xxiii + 897 pp. LCCN QA76.73.C154 D73 2003.

Davis:2005:CD

[Dav05] Stephanie R. Davis. C# 2005 for dummies. John Wiley, New York, NY, USA, second edition, 2005. ISBN 0-7645-9704-3 (paper-

back). ???? pp. LCCN ???? URL http://www.loc.gov/catdir/
enhancements/fy0628/2005927620-d.html.

Dumbill:2004:MDN

[DB04] Edd Dumbill and Niel M. Bornstein. Mono: a developer’s notebook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00792-2. xxii + 278 pp. LCCN QA76.76.M52 D85 2004.

deChamplain:2005:CPG

[dCP05] Michel de Champlain and Brian G. Patrick. C# 2.0: practical guide for programmers. Morgan Kaufmann practical guides series. Elsevier, Amsterdam, The Netherlands, 2005. ISBN 0-12-167451-7. xix + 251 pp. LCCN QA76.73.C154 C45 2005; 06.E01641. URL http://www.

loc.gov/catdir/enhancements/fy0621/2005002332-d.html;
http://www.loc.gov/catdir/toc/ecip057/2005002332.html.

Deitel:2006:CP

[DD06a] Harvey M. Deitel and Paul J. Deitel. C# for programmers. Deitel developer series. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, second edition, 2006. ISBN 0-13-134591-5. xxxvii + 1317 pp. LCCN QA76.73.C154 D455 2006.

Deitel:2006:VCH

[DD06b] Harvey M. Deitel and Paul J. Deitel. Visual C#: how to program. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, second edition, 2006. ISBN 0-13-152523-9. xxxviii + 1591 pp. LCCN QA76.73.C154 D46 2006.

Deitel:2009:CP

[DD09] Harvey M. Deitel and Paul J. Deitel. C# 2008 for programmers. Deitel developer series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 2009. ISBN 0-13-714415-6. xlii + 1251 pp. LCCN QA76.73.C154 D455 2009.

Deitel:2011:VCH

[DD11] Paul J. Deitel and Harvey M. Deitel. Visual C# 2010: how to program. How to program. Pearson Prentice Hall, Upper Saddle River, NJ, USA, fourth edition, 2011. ISBN 0-13-215142-1 (paperback). xxx + 959 pp. LCCN QA76.73.C154 D46 2011.

Deitel:2002:CHP

[Dei02] Harvey M. Deitel. C#: how to program. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-062221-4 (paperback), 0-13-089571-7, 0-13-089572-5, 0-13-064657-1 (CD-ROM). 1568 pp. LCCN QA76.73.C154 C17 2002.

Drossopoulou:2006:FMD

[DLE06] Sophia Drossopoulou, Giovanni Lagorio, and Susan Eisenbach. A flexible model for dynamic linking in Java and C#. Theoretical Computer Science, 368(1–2):1–29, December 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Draheim:2005:GPC

[DLW05] Dirk Draheim, Christof Lutteroth, and Gerald Weber. Generative programming for C#. ACM SIGPLAN Notices, 40(8):29–33, August 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Doyle:2008:CPP

[Doy08] Barbara Doyle. C# programming: from problem analysis to program design. Course Technology, Boston, MA, USA, second edition, 2008. ISBN 1-4239-0146-0 (paperback). xxxi + 1000 pp. LCCN QA76.73.C154 D58 2008.

Doyle:2010:CPP

[Doy10] Barbara Doyle. C# Programming: From Problem Analysis to Program Design. Course Technology, Boston, MA, USA, third edition, 2010. ISBN 0-538-45302-8. 832 (est.) pp. LCCN ???? US$129.95.

Derezinska:2012:QEO

[DR12] Anna Derezińska and Marcin Rudnik. Quality evaluation of objectoriented and standard mutation operators applied to C# programs. Lecture Notes in Computer Science, 7304:42–57, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-642-
30561-0_5/;   http://link.springer.com/content/pdf/10.1007/
978-3-642-30561-0_5.

Davis:2006:CD

[DS06] Stephen R. Davis and Chuck Sphar. C# 2005 for dummies. For dummies. John Wiley, New York, NY, USA, 2006. ISBN 0-7645-9704-3 (paperback). xviii + 406 pp. LCCN 06.E03130; QA76.73.C154 DAV; QA76.73.C154 D34 2006; CD06.S00117. Includes CD-ROM.

Derezinska:2012:OOT

[DS12] Anna Derezińska and Anna Szustek. Object-oriented testing capabilities and performance evaluation of the C# mutation system. Lecture Notes in Computer Science, 7054:229–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28038-2_18/.

Dunn:2003:CAI

[Dun03] Kyle Dunn. A C# application from inspiration to implementation. Wrox Press, Chicago, IL, USA, 2003. ISBN 1-86100-754-X (paperback). vii + 360 pp. LCCN ????

Eaddy:2001:CVJ

[Ead01] Marc Eaddy. C# versus Java. Dr. Dobb’s Journal of Software Tools, 26(2):74, 76, 78, 80, 82, February 2001. CODEN DDJOEB.

ISSN 1044-789X. URL http://www.ddj.com/ftp/2001/2001_02/
csharpj.txt.

ECMA-334-1

[ECM01] ECMA. ECMA-334: C# Language Specification. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2001. ISBN

????  xiii + 479 pp. LCCN ????  URL http://www.

ecma-international.org/publications/files/ecma-st/ECMA-
334.pdf;   http://www.ecma-international.org/publications/
standards/Ecma-334.htm.

ECMA-334-2

[ECM02a] ECMA. ECMA-334: C# Language Specification. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 2002. ISBN ???? xiv + 471 pp. LCCN ???? URL http://www.

ecma-international.org/publications/files/ecma-st/ECMA-
334.pdf;   http://www.ecma-international.org/publications/
standards/Ecma-334.htm.

ECMA:2002:ECL

[ECM02b] ECMA. ECMA-334: C# Language Specification. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 2002. ISBN ???? xiv + 471 pp. LCCN ???? URL http://www.

ecma-international.org/publications/files/ecma-st/ECMA-
334.pdf;   http://www.ecma-international.org/publications/
standards/Ecma-334.htm.

ECMA-334-3

[ECM05] ECMA. ECMA-334: C# Language Specification. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, June 2005. ISBN ???? xvii + 523 pp. LCCN ???? URL http://www.

ecma-international.org/publications/files/ecma-st/ECMA-
334.pdf;   http://www.ecma-international.org/publications/
standards/Ecma-334.htm.

ECMA-334-4

[ECM06] ECMA. ECMA-334: C# Language Specification. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, fourth edition, June 2006. xix + 531

pp. URL http://www.ecma-international.org/publications/
files/ecma-st/ECMA-334.pdf;   http://www.ecma-international.
org/publications/standards/Ecma-334.htm.

Evjen:2010:PAN

[ERH10] Bill Evjen, Devin Rader, and Scott Hanselman. Professional ASP.NET 4 in C# and VB. Wrox Press, Chicago, IL, USA, 2010. ISBN 0-470-50220-7. lviii + 1477 pp. LCCN TK5105.8885.A26 E96 2010.

Erickson:2001:DDE

[Eri01] Jonathan Erickson. Dr. Dobb’s Excellence In Programming Award. Dr. Dobb’s Journal of Software Tools, 26(5):20–21, May 2001. CO-DEN DDJOEB. ISSN 1044-789X. URL http://www.ddj.com/.

Faxen:2008:WWS

[Fax08] Karl-Filip Faxén. Wool — a work stealing library. ACM SIGARCH Computer Architecture News, 36(5):93–100, December 2008. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Fekete:2008:TSD

[Fek08] Alan D. Fekete. Teaching students to develop thread-safe Java classes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):119–123, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Fahnenstich:2006:REP

[FH06] Klaus Fahnenstich and Rainer G.. Haselier. Richtig einsteigen: programmieren lernen mit Visual C# 2005. Richtig einsteigen . . .. Microsoft Press, Unterschleicheim, Germany, 2006. ISBN 3-86645-201-2. 384 pp. LCCN ???? EUR 24.90.

Foxall:2002:CU

[FHCM^{+}02] James Foxall, Wendy Haro-Chun, Ammerins Moss-de Boer, et al. C# in 24 uur. (Dutch) [C# in 24 hours]. In 24 uur; In 24 uur. Sams, Amsterdam, The Netherlands, 2002. ISBN 90-430-0628-9. xix + 529 pp. LCCN ????

Freeman:2003:PNS

[FJ03] Adam Freeman and Allen Jones. Programming .NET security. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00442-7. xviii + 693 pp. LCCN TK5105.59 .F68 2003.

Fraser:2002:BCX

[FL02] Stewart Fraser and Steven Livingstone. Beginning C# XML. Wrox Press, Chicago, IL, USA, 2002. ISBN 1-86100-628-4 (paperback). ix + 729 pp. LCCN M03.E16271. US$28.99.

Ferrara:2008:SUC

[FLF08] Pietro Ferrara, Francesco Logozzo, and Manuel Fahndrich. Safer unsafe code for .NET. ACM SIGPLAN Notices, 43(10):329–346, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ferrara:2002:PNW

[FM02] Alex Ferrara and Matthew MacDonald. Programming .NET Web Services. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00250-5. xvi + 396 pp. LCCN QA76.625 .F49 2002. US$39.95.

URL http://www.oreilly.com/catalog/prognetws.

Fossum:2005:CFC

[Fos05] T. V. Fossum. Classes as first class objects in an environment-passing interpreter. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):261–265, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Foxall:2004:STY

[Fox04] James D. Foxall. Sams teach yourself Microsoft Visual C# .NET 2003 in 24 hours. Sams teach yourself in 24 hours. Howard W. Sams, Indianapolis, IN 46268, USA, second edition, 2004. ISBN 0-672-32538-1, 0-7686-6064-5 (electronic book). xxi + 571 pp. LCCN QA76.73.C154 F694 2004eb. URL http://link.library.

utoronto.ca/eir/EIRdetail.cfm?Resources\_\_ID=367815&T=;
resource.

Foxall:2006:STY

[Fox06] James Foxall. Sams teach yourself Visual C# 2005 in 24 hours: complete starter kit. Howard W. Sams, Indianapolis, IN 46268, USA, 2006. ISBN 0-672-32740-6. ???? pp. LCCN ????

Foxall:2008:STY

[Fox08] James D. Foxall. Sams teach yourself Visual C# 2008 in 24 hours: complete starter kit. Howard W. Sams, Indianapolis, IN 46268, USA, 2008. ISBN 0-672-32990-5. xi + 512 pp. LCCN

QA76.73.C154 F68 2008. URL http://www.loc.gov/catdir/toc/
ecip0819/2008022406.html.

Ferguson:2002:CB

[FPB^{+}02] Jeff Ferguson, Brian Patterson, Jason Beres, Pierre Boutquin, and Meeta Gupta, editors. C# bible. John Wiley, New York, NY, USA, 2002. ISBN 0-7645-4834-4 (paperback). xxxi + 798 pp. LCCN QA76.73.C154 C15 2002.

Fruja:2010:TPT

[Fru10] Nicu G. Fruja. Towards proving type safety of C#. Computer Languages, Systems and Structures, 36(1):60–95, April 2010. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1477842409000128.

Farrell-Vinay:2009:BRM

[FV09] Peter Farrell-Vinay. Book review: Model-based software testing and analysis in C# by Jonathan Jacky, Margus Veanes, Colin Campbell, and Wolfram Schulte, and published by CUP, 2008, (paperback) ISB-10: 0-521-68761-6, 349 pp. ACM SIGSOFT Software Engineering Notes, 34(5):60, September 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fahndrich:2007:EOI

[FX07] Manuel Fahndrich and Songtao Xia. Establishing object invariants with delayed types. ACM SIGPLAN Notices, 42(10):337–350, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gannon:2001:JCC

[G^{+}01] Dennis Gannon et al. Panel discussion: Java, C++, C#, and virtual machines for high performance. In ACM [ACM01], page ?? ISBN 1-58113-359-6. LCCN QA76.9.O35 A26 2001.

Griffiths:2010:PC

[GAL10] Ian Griffiths, Matthew Adams, and Jesse Liberty. Programming C# 4.0. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, sixth edition, 2010. ISBN 0-596-15983-8. xxi + 830 pp. LCCN ????

Grossu:2010:CCC

[GBJ^{+}10] I. V. Grossu, C. Besliu, Al. Jipa, C. C. Bordeianu, D. Felea, E. Stan, and T. Esanu. Code C# for chaos analysis of relativistic many-body systems. Computer Physics Communications, 181(8):1464–1470, August 2010. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465510001384.

Grossu:2012:CCC

[GBJ^{+}12] I. V. Grossu, C. Besliu, Al. Jipa, E. Stan, T. Esanu, D. Felea, and C. C. Bordeianu. Code C# for chaos analysis of relativistic manybody systems with reactions. Computer Physics Communications, 183(4):1055–1059, April 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0010465512000124.

Grossu:2013:CMB

[GBJ^{+}13] I. V. Grossu, C. Besliu, Al. Jipa, D. Felea, T. Esanu, E. Stan, and C. C. Bordeianu. Chaos Many-Body Engine v03: a new version of code C# for chaos analysis of relativistic many-body systems with reactions. Computer Physics Communications, 184(4):1346– 1347, April 2013. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465512004080.

Gatrell:2015:ERC

[GC15] M. Gatrell and S. Counsell. The effect of refactoring on change and fault-proneness in commercial C# software. Science of Computer Programming, 102(??):44–56, May 1, 2015. CODEN SCPGD4.

ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167642314005711.

Goulding:2007:CGD

[GD07] Tom Goulding and Rita DiTrolio. Complex game development by freshman computer science majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):92–99, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927

(electronic). URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.
ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Galassi:2002:GSL

[GDT^{+}02] Mark Galassi, Jim Davies, James Theiler, Brian Gough, Gerard Jungman, Michael Booth, and Fabrice Rossi. GNU Scientific Library: Reference Manual. Network Theory Ltd., Bristol, UK, second edition, 2002. ISBN 0-9541617-3-4. xvi + 601 pp. LCCN QA76.73.C15.

Gerlach:2003:GPS

[GK03] J. Gerlach and J. Kneis. Generic programming for scientific computing in C++, Java, and C#. Lecture Notes in Computer Science, 2834:301–310, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goode:2002:BAN

[GKM^{+}02] Chris Goode, John Kauffman, Christopher L. Miller, Neil Raybould, S. Srinivasa Sivakumar, Dave Sussman, Ollie Cornes, Rob Birdwell, Matt Butler, Gary Johnson, Ajoy Krishnamoorthy, Juan T. Libre, and Chris Ullman. Beginning ASP.NET 1.0 with C#. Programmer to programmer. Wrox Press, Chicago, IL, USA, 2002. ISBN 1-86100-734-5. xiii + 800 pp. LCCN ????

Goode:2003:BAN

[GKM^{+}03] Chris Goode, John Kauffman, Christopher Miller, Neil Raybould, Srinivasa Sivakumar, David Sussman, Ollie Cornes, Rob Birdwell, Matt Butler, Gary Johnson, Ajoy Krishnamoorthy, Juan T. Libre, and Chris Ullman. Beginning ASP.NET 1.0 with C#. Programmer to programmer. Wrox Press, Chicago, IL, USA, 2003. ISBN 0-7645-4370-9 (paperback). xiii + 848 pp. LCCN TK5105.8885.A26

B44 2003b. US$28.99. URL http://www.loc.gov/catdir/bios/
wiley047/2003107065.html.

Garcia:2021:IHG

[GMN21] Andrés Amaya García, David May, and Ed Nutting. Integrated hardware garbage collection. ACM Transactions on Embedded Computing Systems, 20(5):40:1–40:25, July 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://dl.acm.org/doi/10.

1145/3450147.

Gunnerson:2001:PCI

[GN01] Eric Gunnerson and Wout Natter. Programmeren in C#: een inleiding voor programmeurs. (Dutch). [Programming in C#: an introduction for programmers]. Academic Service, Schoonhoven, The Netherlands, 2001. ISBN 90-395-1717-7. xxi + 338 pp. LCCN ????

Goulding:2008:CGD

[Gou08] Tom Goulding. Complex game development throughout the college curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):68–71, December 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gordon:2012:URI

[GPP^{+}12] Colin S. Gordon, Matthew J. Parkinson, Jared Parsons, Aleks Bromfield, and Joe Duffy. Uniqueness and reference immutability for safe parallelism. ACM SIGPLAN Notices, 47(10):21–40, October 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grinzo:2002:PBC

[Gri02] Lou Grinzo. Programmer’s bookshelf: C# and the .NET platform. Dr. Dobb’s Journal of Software Tools, 27(1):108, January 2002. CO-DEN DDJOEB. ISSN 1044-789X. URL http://www.ddj.com/.

Griffiths:2012:PC

[Gri12a] Ian Griffiths. Programming C# 5.0. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-2041-4. 886 pp. LCCN ????

Griffiths:2012:PCB

[Gri12b] Ian Griffiths. Programming C# 5.0: Building Windows 8 Metro, Web, and Desktop Applications for the .NET 4.5 Framework. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-2041-4 (paperback). 800 pp. LCCN ???? EUR 45.00, UK £42.50.

Grossu:2022:MHF

[Gro22] I. V. Grossu. Migration of hyper-fractal analysis from Visual Basic 6 to C# .Net. Computer Physics Communications, 271(??): Article 108189, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0010465521003015.

Geller:2005:TME

[GS05] Viktor Geller and Christelle Scharff. Traditional and more “exotic” .NET languages: VB .NET, J#, C# and SML .NET. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):406, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gunnerson:2005:PIC

[GWH05] Eric Gunnerson, Nick Wienholt, and Anders Hejlsberg. A programmer’s introduction to C# 2.0. Apress, Berkeley, CA, USA, third edition, 2005. ISBN 1-59059-501-7 (paperback). xxxvii + 529 pp. LCCN QA76.73.C153 G86 2005. URL http://link.library.

utoronto.ca/eir/EIRdetail.cfm?Resources\_\_ID=362636&T=;
resource.

Harvey:2000:CPP

[H^{+}00] Burton Harvey et al., editors. C# programming with the public beta. Programmer to programmer. Wrox Press, Chicago, IL, USA, 2000. ISBN 1-86100-487-7 (paperback). ix + 393 pp. LCCN QA76.73.C154 C173 2000.

Hertz:2006:GOL

[HBM^{+}06] Matthew Hertz, Stephen M. Blackburn, J. Eliot B. Moss, Kathryn S. McKinley, and Darko Stefanović. Generating object lifetime traces with Merlin. ACM Transactions on Programming Languages and Systems, 28(3):476–516, May 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Herbert:2002:CCR

[Her02] Schildt Herbert. C#: the complete reference. Osborne/McGraw-Hill, Berkeley, CA, USA, 2002. ISBN 0-07-213485-2. xxiv + 933 pp. LCCN QA76.73.C154 S34 2002b.

Hashmi:2012:CNI

[HNTL12] Atif Hashmi, Andrew Nere, James Jamal Thomas, and Mikko Lipasti. A case for neuromorphic ISAs. ACM SIGPLAN Notices, 47 (4):145–158, April 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hoffman:2003:ANC

[Hof03] Kevin Hoffman. ASP.NET e-commerce programming: problem, design, solution. Programmer to programmer. Wrox Press, Chicago, IL, USA, [c#] edition, 2003. ISBN 1-86100-803-1. ix + 342 pp. LCCN BLi 991861008031.

Hoffman:2006:MVC

[Hof06] Kevin Hoffman. Microsoft Visual C# 2005 unleashed. Howard W. Sams, Indianapolis, IN 46268, USA, 2006. ISBN 0-672-32776-7 (paperback). xix + 692 pp. LCCN ????

Holzner:2004:MVC

[Hol04] Steven Holzner. Microsoft Visual C# .NET 2003 kick start. Howard W. Sams, Indianapolis, IN 46268, USA, 2004. ISBN 0-672-32547-0. xiii + 604 pp. LCCN QA76.73.C154 H65 2003.

Hanson:2004:RCC

[HP04] David R. Hanson and Todd A. Proebsting. A research C# compiler. Software—Practice and Experience, 34(13):1211–1224, November 10, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-

024X (electronic). URL ftp://ftp.research.microsoft.com/
pub/tr/tr-2003-32.pdf;   http://research.microsoft.com/
research/pubs/view.aspx?msr_tr_id=MSR-TR-2003-32.

Hawblitzel:2009:AVP

[HP09] Chris Hawblitzel and Erez Petrank. Automated verification of practical garbage collectors. ACM SIGPLAN Notices, 44(1):441–453, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Huddleston:2005:BCD

[HRA^{+}05] James Huddleston, Ranga Raghuram, Scott Allen, Syed Fahad Gilani, Jacob Hammer Pedersen, and Jon Reid, editors. Beginning C# databases: from novice to professional. Apress, Berkeley, CA, USA, 2005. ISBN 1-59059-433-9. xxiv + 622 pp. LCCN

QA76.73.C154 B445 2005. URL http://www.columbia.edu/cgi-
bin/cul/resolve?clio5291165.

Huang:2008:ESS

[HS08] Shan Shan Huang and Yannis Smaragdakis. Expressive and safe static reflection with MorphJ. ACM SIGPLAN Notices, 43(6):79–89, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Homer:2003:FLA

[HSF03] Alex Homer, David Sussman, and Mark Fussell. A first look at ADO.NET and System.XML v.2.0. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-321-22839-1. xxi + 297 pp. LCCN QA76.625 .H655 2003.

Homer:2004:FLA

[HSH04] Alex Homer, David Sussman, and Rob Howard. A first look at ASP.NET v. 2.0. Microsoft .NET development series. Addison-Wes-ley, Reading, MA, USA, 2004. ISBN 0-321-22896-0 (paperback). xxviii + 498 pp. LCCN TK5105.8885.A26 H6598 2004.

Hilyard:2006:CC

[HT06] Jay Hilyard and Stephen Teilhet. C# cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-10063-9. xxviii + 1154 pp. LCCN QA76.73.C154 T43 2004.

Hilyard:2007:CC

[HT07] Jay Hilyard and Stephen Teilhet. C# 3.0 cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2007. ISBN 0-596-51610-X (paperback). xxvi + 857 pp. LCCN QA76.73.C154; QA76.73.C154eb.

URL http://www.oreilly.com/catalog/9780596516109.

Hejlsberg:2008:CPL

[HTWG08] Anders Hejlsberg, Mads Togersen, Scott Wiltamuth, and Peter Golde, editors. The C# Programming Language. Addison-Wesley, Reading, MA, USA, third edition, 2008. ISBN 0-321-56299-2. xviii + 754 pp. LCCN QA76.73.C154 H45 2008.

Hejlsberg:2011:CPL

[HTWG11] Anders Hejlsberg, Mads Togersen, Scott Wiltamuth, and Peter Golde, editors. The C# Programming Language. Addison-Wesley, Reading, MA, USA, fourth edition, 2011. ISBN 0-321-74176-5. xviii + 844 pp. LCCN QA76.73.C154. US$44.99.

URL http://www.pearsonhighered.com/bookseller/product/
C-Programming-Language-Covering-C-40-The/9780321741769.
page.

Humphrey:2003:GCC

[Hum03] William F. Humphrey. Generalized callbacks: C++ and C#. Dr. Dobb’s Journal of Software Tools, 28(3):42–43, 46–47, March 2003. CODEN DDJOEB. ISSN 1044-789X. URL http://www.ddj.com/

documents/s=7826/ddj0303e/.

Hummel:2006:LFD

[Hum06] Joe Hummel. LINQ, the future of data access in C# 3.0. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52841-8. 64 pp. LCCN QA76.73.C154; QA76.73.C154 H86 2006eb. URL http:

//www.oreilly.com/catalog/9780596528416.

Hunt:2003:LSM

[Hun03] J. Hunt. Look sharp! Microsoft’s C# column has often been described as a Java killer, and the languages have a lot in common. Application Development Advisor, 7(2):32–35, 2003. CODEN ???? ISSN 1369-4200.

Hejlsberg:2004:CPL

[HWG04] Anders Hejlsberg, Scott Wiltamuth, and Peter Golde. The C# programming language. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-321-15491-6 (hardcover). xiv + 644 pp. LCCN QA76.76.C154 H45 2004.

Hejlsberg:2006:CPL

[HWG06] Anders Hejlsberg, Scott Wiltamuth, and Peter Golde. The C# programming language. Addison-Wesley, Reading, MA, USA, second edition, 2006. ISBN 0-321-33443-4 (hardback). xiv + 704 pp. LCCN

QA76.73.C154 H45 2006. URL http://www.loc.gov/catdir/toc/
ecip0613/2006015047.html.

ISO:2003:IIIb

[Int03] International Organization for Standardization. ISO/IEC 23270:2003: Information technology — C# Language Specification. International Organization for Standardization, Geneva, Switzerland, 2003. ISBN ???? xiii + 471 pp. LCCN ???? 328 Sfr (print). URL http://

standards.iso.org/ittf/PubliclyAvailableStandards/c036768_
ISO_IEC_23270_2003(E).zip;   http://www.iso.org/iso/en/
CatalogueDetailPage.CatalogueDetail?CSNUMBER=36768.

ISO:2005:IIIa

[Int05a] International Organization for Standardization. ISO/IEC 23271:2003: Information technology — Common Language Infrastructure. International Organization for Standardization, Geneva, Switzerland, 2005. ISBN ???? xi + 99 (Part. I), ix + 164 (Part. II), vi + 125 (Part. III), iii + 16 (Part. IV), iv + 79 (Part. V) pp. LCCN

???? 328 Sfr (print). URL http://standards.iso.org/ittf/
PubliclyAvailableStandards/c036769_ISO_IEC_23271_2003(E)
.zip;   http://www.iso.org/iso/en/CatalogueDetailPage.CatalogueDetail?
CSNUMBER=36769.

ISO:2005:IIIb

[Int05b] International Organization for Standardization. ISO/IEC TR 23272: Information technology — Common Language Infrastructure — Profiles and Libraries. International Organization for Standardization, Geneva, Switzerland, 2005. ISBN ???? 6 pp. LCCN ???? 328 Sfr (print). URL http://standards.iso.

org/ittf/PubliclyAvailableStandards/c036770_ISO_IEC_TR_
23272_2003(E).zip;   http://www.iso.org/iso/en/CatalogueDetailPage.
CatalogueDetail?CSNUMBER=36770.

ISO:2006:IIIb

[Int06] International Organization for Standardization. ISO/IEC 23270:2006: Information technology — Programming languages — C#. Technical report. International Organization for Standardization, Geneva,

Switzerland, 2006. URL http://www.iso.org/iso/iso_catalogue/
catalogue_tc/catalogue_detail.htm?csnumber=42926.

Jones:2010:VCR

[JF10] Allen Jones and Adam Freeman, editors. Visual C# 2010 recipes: a problem-solution approach. Apress, Berkeley, CA, USA, 2010. ISBN 1-4302-2525-4. xxiii + 987 pp. LCCN ????

Jimeno:2007:CYO

[JJOA07] Roberto Jimeno, M. S. Jorge, and L. Ortega-Arjona. Curb your objects!: an orthodox form for C# classes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):138– 141, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-

3927 (electronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Jantz:2013:ESM

[JK13] Michael R. Jantz and Prasad A. Kulkarni. Exploring single and multilevel JIT compilation policy for modern machines 1. ACM Transactions on Architecture and Code Optimization, 10(4):22:1–22:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jagannathan:2014:ARV

[JLP^{+}14] Suresh Jagannathan, Vincent Laporte, Gustavo Petri, David Pichardie, and Jan Vitek. Atomicity refinement for verified compilation. ACM Transactions on Programming Languages and Systems, 36(2):6:1–6:??, July 2014. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Jones:2006:VCR

[JM06] Allen Jones and Matthew MacDonald. Visual C# 2005 Recipes: a Problem-Solution Approach. Apress, Berkeley, CA, USA, 2006. ISBN 1-59059-589-0. xxi + 565 pp. LCCN ????

Jones:2002:STY

[Jon02] Bradley Jones. Sams teach yourself C# 21 dager. IDG Norge Books, Oslo, Norway, 2002. ISBN 82-7772-242-7. x + 735 pp. LCCN ????

Jones:2004:STY

[Jon04] Bradley Jones. Sams teach yourself the C# language in 21 days. Sams teach yourself in 21 days. Howard W. Sams, Indianapolis, IN 46268, USA, third edition, 2004. ISBN 0-672-32546-2 (paper), 0-7686-6058-0 (electronic). xxi + 779 pp. LCCN QA76.73.C154

J67 2004eb.  URL http://link.library.utoronto.ca/eir/
EIRdetail.cfm?Resources\_\_ID=367816 and T=; resource.

Jagger:2007:ACS

[JPS07] Jon Jagger, Nigel Perry, and Peter Sestoft. Annotated C# standard. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2007. ISBN 0-12-372511-9. xxiii + 825 pp. LCCN QA76.73.C154 J35 2007.

Jacobs:2008:PMC

[JPS^{+}08] Bart Jacobs, Frank Piessens, Jan Smans, K. Rustan M. Leino, and Wolfram Schulte. A programming model for concurrent objectoriented programs. ACM Transactions on Programming Languages and Systems, 31(1):1:1–1:48, December 2008. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Kiselyov:2017:SFC

[KBPS17] Oleg Kiselyov, Aggelos Biboudis, Nick Palladinos, and Yannis Smaragdakis. Stream fusion, to completeness. ACM SIGPLAN Notices, 52(1):285–299, January 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kedia:2017:SFS

[KCP^{+}17] Piyus Kedia, Manuel Costa, Matthew Parkinson, Kapil Vaswani, Dimitrios Vytiniotis, and Aaron Blankstein. Simple, fast, and safe manual memory management. ACM SIGPLAN Notices, 52(6):233– 247, June 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kent:2005:VCD

[Ken05] Jeff Kent. Visual C# 2005 Demystified: a Self-Teaching Guide. Mc-Graw-Hill, New York, NY, USA, 2005. ISBN 0-07-226170-6. xxii + 352 pp. LCCN ????

Kiczales:2004:CLG

[Kic04] G. Kiczales. Crosscut: A little goes a long way: Attributes in C# and annotations in Java provide a hook for principled macros and metaprogramming in the C family. Lisp has had this technology for years, offering many lessons. Software Development, 12(5):52–53, 2004. CODEN ???? ISSN 1070-8588.

Kimmel:2002:ACP

[Kim02] Paul Kimmel. Advanced C# programming. Osborne/McGraw-Hill, Berkeley, CA, USA, 2002. ISBN 0-07-222417-7 (paperback). xxiii + 551 pp. LCCN QA76.73.C154 K56 2002eb. URL http://www.

columbia.edu/cgi-bin/cul/resolve?clio5385247.

Kauffman:2002:BAN

[KMM^{+}02] John Kauffman, Brian Matsik, Eric N. Mintz, Jan D. Narkiewicz, Kent Tegels, John West, Donald Xie, Jesudas Chinnathampi, Fabio Claudio Ferracchiati, and James Greenwood. Beginning ASP.NET databases using C#. Programmer to programmer. Wrox Press, Chicago, IL, USA, 2002. ISBN 1-86100-741-8. vii + 453 pp. LCCN ????

Kalibera:2011:SRT

[KPHV11] Tomas Kalibera, Filip Pizlo, Antony L. Hosking, and Jan Vitek. Scheduling real-time garbage collection on uniprocessors. ACM Transactions on Computer Systems, 29(3):8:1–8:??, August 2011. CODEN ACSYEC. ISSN 0734-2071.

Kaiser:2006:CJC

[KPPÉR06] Claude Kaiser, Jean-François Pradat-Peyre, Sami Évangelista, and Pierre Rousseau. Comparing Java, C# and Ada monitors queuing policies: a case study and its Ada refinement. ACM SIGADA Ada Letters, 26(2):23–37, August 2006. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kliot:2009:LFC

[KPS09] Gabriel Kliot, Erez Petrank, and Bjarne Steensgaard. A lock-free, concurrent, and incremental stack scanning mechanism for garbage collectors. Operating Systems Review, 43(3):3–13, July 2009. CO-DEN OSRED8. ISSN 0163-5980.

Kennedy:2018:TGC

[KR18] Andrew J. Kennedy and Claudio V. Russo. Transposing G to C^{♯}: Expressivity of generalized algebraic data types in an objectoriented language. Theoretical Computer Science, 741(??):1–2, ???? 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0304397518301270.

Kanig:2012:HLC

[KSD12] Johannes Kanig, Edmond Schonberg, and Claire Dross. Hi-Lite: the convergence of compiler technology and program verification. ACM SIGADA Ada Letters, 32(3):27–34, December 2012. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Lauwens:2017:PRC

[Lau17] Ben Lauwens. ResumableFunctions: C# sharp style generators for Julia. Journal of Open Source Software, 2(18):400:1–400:2, October 2017. CODEN ???? ISSN 2475-9066. URL http://joss.theoj.

org/papers/10.21105/joss.00400.

Lerman:2010:PEF

[Ler10] Julia Lerman. Programming Entity Framework. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2010. ISBN 0-596-80727-9, 0-596-80726-0. ???? pp. LCCN ????

Lewis:2006:CSS

[Lew06] John Lewis. C# software solutions: foundations of program design. Pearson/Addison-Wesley, Boston, MA, USA, 2006. ISBN 0-321-26716-8, 0-321-41720-8 (International ed.). xxi + 712 pp. LCCN

QA76.73.C154 L38 2006. URL http://www.loc.gov/catdir/toc/
ecip064/2005035820.html.

Liberty:2002:PAN

[LH02] Jesse Liberty and Dan Hurwitz. Programming ASP.NET. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00171-1. xiii + 944 pp. LCCN QA76.73.C154 L52 2002; TK5105.8885.A26 L53

2002. US$49.95. URL http://safari.oreilly.com/0596001711;
 http://www.oreilly.com/catalog/progaspdotnet.

Liberty:2003:PAN

[LH03] Jesse Liberty and Dan Hurwitz. Programming ASP.NET. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00487-7. xiii + 988 pp. LCCN TK5105.8885.A26 L53 2003.

Liberty:2004:PNW

[LH04] Jesse Liberty and Dan Hurwitz. Programming .NET Windows applications. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00321-8. xv + 1228 pp. LCCN QA76.76.A65 L52 2004.

Lhotka:2006:ECB

[Lho06] Rockford Lhotka. Expert C# 2005 Business Objects. Apress, Berkeley, CA, USA, second edition, 2006. ISBN 1-59059-632-3. 696 (est.)

pp. LCCN ???? US$59.99. URL http://www.apress.com/book/
bookDisplay.html?bID=10091.

Liberty:2001:PC

[Lib01] Jesse Liberty. Programming C#. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 2001. ISBN 0-596-00117-7. xix + 658 pp. LCCN QA76.73.C154 L432 2001. US$39.95. URL http://www.oreilly.

com/catalog/progcsharp.

Liberty:2002:LC

[Lib02a] Jesse Liberty. Learning C#. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00376-5. xi + 354 pp. LCCN QA76.73.C154 L43

2002. US$34.95. URL http://safari.oreilly.com/0596003765;
 http://www.oreilly.com/catalog/learncsharp.

Liberty:2002:PC

[Lib02b] Jesse Liberty. Programming C#. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00309-9. xviii + 629 pp. LCCN QA76.73.C154 L53 2002. US$39.95.

URL http://safari.oreilly.com/0596003099;   http://www.
oreilly.com/catalog/progcsharp2.

Liberty:2003:PC

[Lib03] Jesse Liberty. Programming C#. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2003. ISBN 0-596-00489-3. xviii + 689 pp. LCCN QA76.76.C154 L53 2003.

Liberty:2005:PC

[Lib05a] Jesse Liberty. Programming C#. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2005. ISBN 0-596-00699-3. xx + 644 pp. LCCN QA76.73.C154 L53 2005. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/( );   http://proquest.
safaribooksonline.com/0596006993.

Liberty:2005:VCD

[Lib05b] Jesse Liberty. Visual C# 2005: a developer’s notebook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00799-X (paperback). xv + 221 pp. LCCN QA76.73.C154 L435 2005. URL

http://proquest.safaribooksonline.com/059600799X.

Lippman:2001:CD

[Lip01] Stanley B. Lippman. The C# delegate. C/C++ Users Journal, 19 (7):42–??, July 2001. CODEN CCUJEX. ISSN 1075-2838.

Liu:2006:II

[LKM06] Jed Liu, Aaron Kimball, and Andrew C. Myers. Interruptible iterators. ACM SIGPLAN Notices, 41(1):283–294, January 2006. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lutz:2003:CNF

[LL03a] Michael H. Lutz and Phillip A. Laplante. C# and the .NET framework: Ready for real time? IEEE Software, 20(1):74–80, January/ February 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-

7459 (electronic). URL http://dlib.computer.org/so/books/
so2003/pdf/s1074.pdf;   http://www.computer.org/software/
so2003/s1074abs.htm.

Lutz:2003:ISR

[LL03b] Michael H. Lutz and Phillip A. Laplante. IEEE Software: Realtime systems: C# and the .NET framework: Ready for real time? IEEE Distributed Systems Online, 4(2), 2003. ISSN 1541-4922. URL

http://dsonline.computer.org/0302/f/sp1lap.htm.

Liberty:2006:LC

[LM06] Jesse Liberty and Brian MacDonald. Learning C# 2005: Get Started with C# 2005 and .NET Programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 1-59059-589-0, 0-596-10209-7. 544 (est.) pp. LCCN ????

Liberty:2008:LC

[LM08] Jesse Liberty and Brian MacDonald. Learning C#. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-15501-8. 704 (est.) pp. LCCN QA76.73.C153 L53

Lee:2004:JBN

[LN04] Y. Lee and S. Na. Java bytecode-to-.NET MSIL translator for construction of platform independent information systems. Lecture Notes in Computer Science, 3215:826–832, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lowy:2010:PWS

[Löw10] Juval Löwy. Programming WCF services. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2010. ISBN 1-4493-8247-9, 0-596-80548-9. ???? pp. LCCN ????

Lehoczky:2018:HLN

[LRT^{+}18] Zoltán Lehóczky, András Retzler, Richárd Tóth, Álmos Szabó, Benedek Farkas, and Krisztián Somogyi. High-level .NET software implementations of unum Type I and posit with simultaneous FPGA implementation using Hastlayer. In ACM [ACM18], pages 1–7.

Liu:2014:FFL

[LTZ14] Peng Liu, Omer Tripp, and Xiangyu Zhang. Flint: fixing linearizability violations. ACM SIGPLAN Notices, 49(10):543–560, October 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lutz:2003:BBC

[Lut03] Michael J. Lutz. Bookshelf: Building C# programs with patterns [C# Design Patterns]; using Star Schema in Oracle [Oracle DBA Guide to Data Warehousing and Star Schemas]; immersive Java text [Java Outside In]; applying AI to the Web [Web Intelligence]; moronic marketing [In Search of Stupidity: Over 20 Years of High-Tech Marketing Disasters]. Computer, 36(9):96–97, September 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/dl/mags/co/2003/09/r9096.
htm;   http://csdl.computer.org/dl/mags/co/2003/09/r9096.
pdf.

Liberty:2008:PC

[LX08] Jesse Liberty and Donald Xie. Programming C# 3.0. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fifth edition, 2008. ISBN 0-596-52743-8. xviii + 587 pp. LCCN QA76.73.C154. URL http://www.

oreilly.com/catalog/9780596527433.

Makofske:2003:MSP

[MA03] David Makofske and Kevin Almeroth. Multicast sockets: practical guide for programmers. The Morgan Kaufmann practical guides series. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-846-X. xii + 180 pp. LCCN TK5105.887 .M35 2003. US$19.95. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/;   http://www.loc.gov/
catdir/description/els051/2004555479.html;   http://www.
loc.gov/catdir/toc/els051/2004555479.html.

MacBeth:2004:CPH

[Mac04] Gregory S. MacBeth. C# programmer’s handbook. Apress, Berkeley, CA, USA, 2004. ISBN 1-59059-270-0. xix + 561 pp. LCCN QA76.73.C154 M3344 2004b.

MacDonald:2005:BAN

[Mac05a] Matthew MacDonald. Beginning ASP.NET 1.1 in C#: from novice to professional. Apress, Berkeley, CA, USA, 2005. ISBN 1-59059-431-2. xxii + 1017 pp. LCCN TK5105.8885.A26 M33 2005b. URL

http://www.columbia.edu/cgi-bin/cul/resolve?clio5459160.

MacDonald:2005:PAN

[Mac05b] Matthew MacDonald, editor. Pro ASP.NET 1.1 in C#: from professional to expert. Apress, Berkeley, CA, USA, 2005. ISBN 1-59059-351-0. LCCN TK5105.8885.A26 P76 2005b; **See. URL http://

library.books24x7.com/library.asp?B&isbn=1590593510.

Matthew:2006:PNW

[Mac06] Matthew MacDonald. Pro.NET 2.0 Windows forms and custom controls in C#. Apress, Berkeley, CA, USA, 2006. ISBN 1-59059-439-8. xxxix + 1037 pp. LCCN ????

MacDonald:2008:PWC

[Mac08] Matthew MacDonald. Pro WPF in C# 2008: Windows Presentation Foundation with .NET 3.5. Apress, Berkeley, CA, USA, second edition, 2008. ISBN 1-59059-955-1. xxix + 1040 pp. LCCN ????

MacDonald:2010:WC

[Mac10] Matthew MacDonald. WPF in C# 2010. Apress, Berkeley, CA, USA, third edition, 2010. ISBN 1-4302-7205-8 (paperback). LCCN ???? US$54.99.

Magennis:2010:LOU

[Mag10] Troy Magennis. LINQ to objects using C# 4.0: using and extending LINQ to objects and Parallel LINQ (PLINQ). The Addison-Wesley Microsoft technology series. Addison-Wesley, Reading, MA, USA, 2010. ISBN 0-321-63720-8, 0-321-63700-3. xx + 312 pp. LCCN ????

Marshall:2005:PMV

[Mar05] Donis Marshall. Programming Microsoft Visual C# 2005: The Language. Microsoft Press, Redmond, WA, USA, 2005. ISBN 0-7356-2181-0. ???? pp. LCCN ????

Mayo:2008:CNF

[May08] Joseph Mayo. C# 3.0: with the .NET Framework 3.5 unleashed. Howard W. Sams, Indianapolis, IN 46268, USA, second edition, 2008. ISBN 0-672-32981-6. xxiv + 1022 pp. LCCN QA76.73.C154

M38 2008. URL http://www.loc.gov/catdir/toc/ecip0820/
2008026117.html.

McCluskey:2003:CT

[McC03a] Glen McCluskey. C# types. ;login: the USENIX Association newsletter, 28(3):??, June 2003. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.usenix.org/publications/login/2003-
06/pdfs/mccluskey.pdf.

McCluskey:2003:ECH

[McC03b] Glen McCluskey. Examining the C# “hello world” program. ;login: the USENIX Association newsletter, 28(2):??, April 2003. CO-DEN LOGNEM. ISSN 1044-6397. URL http://www.usenix.org/

publications/login/2003-04/pdfs/mccluskey.pdf.

McCluskey:2003:IC

[McC03c] Glen McCluskey. An introduction to C#. ;login: the USENIX Association newsletter, 28(1):??, February 2003. CODEN LOGNEM.

ISSN 1044-6397. URL http://www.usenix.org/publications/
login/2003-02/pdfs/mccluskey.pdf.

McCluskey:2003:UCP

[McC03d] Glen McCluskey. Using C# properties and static members. ;login: the USENIX Association newsletter, 28(5):??, October 2003. CODEN LOGNEM. ISSN 1044-6397. URL http://www.usenix.

org/publications/login/2003-10/pdfs/mccluskey.pdf.

McCluskey:2003:WCC

[McC03e] Glen McCluskey. Working with C# classes. ;login: the USENIX Association newsletter, 28(4):??, August 2003. CODEN LOGNEM.

ISSN 1044-6397. URL http://www.usenix.org/publications/
login/2003-08/pdfs/mccluskey.pdf.

McCluskey:2004:COO

[McC04a] Glen McCluskey. C# overloaded operators. ;login: the USENIX Association newsletter, 29(3):??, June 2004. CODEN LOGNEM. ISSN

1044-6397. URL http://www.usenix.org/publications/login/
2004-06/pdfs/mccluskey.pdf.

McCluskey:2004:MUC

[McC04b] Glen McCluskey. Making use of C# collections. ;login: the USENIX Association newsletter, 29(5):??, October 2004. CODEN LOGNEM.

ISSN 1044-6397. URL http://www.usenix.org/publications/
login/2004-10/pdfs/mccluskey.pdf.

McCluskey:2004:UCA

[McC04c] Glen McCluskey. Using C# abstract classes. ;login: the USENIX Association newsletter, 29(2):??, April 2004. CODEN LOGNEM. ISSN

1044-6397. URL http://www.usenix.org/publications/login/
2004-04/pdfs/McCluskey.pdf.

McCluskey:2004:UCR

[McC04d] Glen McCluskey. Using C# reflection. ;login: the USENIX Association newsletter, 29(4):??, August 2004. CODEN LOGNEM. ISSN

1044-6397. URL http://www.usenix.org/publications/login/
2004-08/pdfs/mccluskey.pdf.

McCluskey:2004:WCI

[McC04e] Glen McCluskey. Working with C# interfaces. ;login: the USENIX Association newsletter, 29(1):??, February 2004. CO-DEN LOGNEM. ISSN 1044-6397. URL http://www.usenix.org/

publications/login/2004-02/pdfs/mccluskey.pdf.

McCluskey:2005:WCS

[McC05] Glen McCluskey. Working with C# serialization. ;login: the USENIX Association newsletter, 30(1):??, February 2005. CO-DEN LOGNEM. ISSN 1044-6397. URL http://www.usenix.org/

publications/login/2005-02/pdfs/mccluskey.pdf.

McKinley:2016:PWU

[McK16] Kathryn S. McKinley. Programming the world of uncertain things (keynote). ACM SIGPLAN Notices, 51(1):1–2, January 2016. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McSherry:2010:PIQ

[McS10] Frank McSherry. Privacy integrated queries: an extensible platform for privacy-preserving data analysis. Communications of the Association for Computing Machinery, 53(9):89–97, September 2010. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Metsker:2004:DPC

[Met04] Steven John Metsker. Design patterns in C#. The software patterns series. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-321-12697-1. xvii + 456 pp. LCCN QA76.73.C154 M48 2004.

Meyer:2002:ACT

[Mey02] Bertrand Meyer. Assessing a C# text. Computer, 35(4):86–88, April 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.org/co/books/co2002/pdf/
r4086.pdf;   http://www.computer.org/computer/co2002/r4086abs.
htm.

Menon:2006:VSP

[MGM^{+}06] Vijay S. Menon, Neal Glew, Brian R. Murphy, Andrew McCreight, Tatiana Shpeisman, Ali-Reza Adl-Tabatabai, and Leaf Petersen. A verifiable SSA program representation for aggressive compiler optimization. ACM SIGPLAN Notices, 41(1):397–408, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Michaelis:2007:EC

[Mic07] Mark Michaelis. Essential C# 2.0. Microsoft .NET development series. Addison-Wesley, Reading, MA, USA, 2007. ISBN 0-321-15077-5 (paperback). xli + 720 pp. LCCN QA76.73.C154 M523 2007. URL

http://www.loc.gov/catdir/toc/ecip0611/2006009699.html.

Michaelis:2009:ECN

[Mic09] Mark Michaelis. Essential C# 3.0: for .NET Framework 3.5. Microsoft .NET development series. Addison-Wesley, Reading, MA, USA, 2009. ISBN 0-321-53392-5 (paperback). xlv + 826 pp. LCCN QA76.73.C154 M5235 2009.

Michaelis:2010:ECH

[Mic10] Mark Michaelis. Essential C# 4.0. Addison-Wesley, Reading, MA, USA, third edition, 2010. ISBN 0-321-69563-1, 0-321-69469-4. li + 930 pp. LCCN QA76.73.C154 M5237 2010.

MacDonald:2003:CDS

[MJ03] Matthew MacDonald and Erik Johansson. C# data security handbook. Wrox Press, Chicago, IL, USA, 2003. ISBN 1-86100-801-5 (paperback). vii + 356 pp. LCCN ????

Murach:2004:MC

[ML04] Joel Murach and Doug Lowe. Murach’s C#. .Net developer’s guide. Mike Murach and Associates, Fresno, CA, USA, 2004. ISBN 1-890774-22-7. xvi + 750 pp. LCCN QA76.73.C154 M86 2004.

Mojica:2002:CVN

[Moj02] Jose Mojica. C# & VB.NET Conversion Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00319-6. iv + 139 pp. LCCN QA76.73.C154 M65 2002. US$14.95.

URL http://safari.oreilly.com/0596003196;   http://www.
oreilly.com/catalog/csharpvbpr.

Mossenbock:2004:CP

[Mös04] Hanspeter Mössenböck. C# to the point. Pearson Education, Harlow, UK, 2004. ISBN 0-321-25290-X (paperback). xiii + 246 pp. LCCN QA76.73.C154 M6713 2005. UK£29.99. Translated from the German original, Softwareentwicklung mit C#, by Bedford Translation Ltd with assistance from David Lightfoot (Oxford Brookes University).

Mossenbock:2006:SCK

[Mös06] Hanspeter Mössenböck. Softwareentwicklung mit C# 2.0: ein kompakter Lehrgang. dpunkt.lehrbuch. dpunkt-Verlag, Heidelberg, Germany, second edition, 2006. ISBN 3-89864-406-5. xiii + 272 pp.

LCCN ???? EUR 29.00. URL http://deposit.ddb.de/cgi-bin/
dokserv?id=2791467&prov=M&dok_var=1&dok_ext=htm; http://
www.gbv.de/dms/ilmenau/toc/508849624moess.PDF.

Montanari:2018:OCC

[MP18] Mattia Montanari and Nik Petrinic. OpenGJK for C, C# and Matlab: Reliable solutions to distance queries between convex bodies in three-dimensional space. SoftwareX, 7(??):352–355, January/ June 2018. CODEN ???? ISSN 2352-7110. URL http://www.

sciencedirect.com/science/article/pii/S2352711018300591.

Moors:2008:GHK

[MPO08] Adriaan Moors, Frank Piessens, and Martin Odersky. Generics of a higher kind. ACM SIGPLAN Notices, 43(10):423–438, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Miller:2004:CLI

[MR04] James S. Miller and Susann Ragsdale. The Common Language Infrastructure Annotated Standard. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-321-15493-2. xxxii + 891 pp. LCCN QA76.7 .M52 2003.

Miura:2009:AGI

[MSK09] Motoki Miura, Taro Sugihara, and Susumu Kunifuji. Anchor Garden: an interactive workbench for basic data concept learning in object oriented programming languages. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):141– 145, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

McCreight:2007:GFC

[MSLL07] Andrew McCreight, Zhong Shao, Chunxiao Lin, and Long Li. A general framework for certifying garbage collectors and their mutators. ACM SIGPLAN Notices, 42(6):468–479, June 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Murach:2006:MC

[Mur06] Joel Murach. Murach’s C# 2005. Mike Murach and Associates, Fresno, CA, USA, 2006. ISBN 1-890774-37-5. 812 (est.) pp.

LCCN ???? US$52.50. URL http://www.murach.com/books/
csh5/index.htm.

Naugler:2008:RPL

[Nau08a] David Naugler. Review of Pro LINQ Object Relational Mapping in C# 2008 by Vijay Mehta, Apress, 2008, $49.99, ISBN: 978-1-59059-965-5. ACM Queue: Tomorrow’s Computing Today, 6(7):41, November/December 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Naugler:2008:RPW

[Nau08b] David Naugler. Review of Pro WPF in C# 2008: Windows Presentation Foundation with .NET 3.5, (2nd ed.),’ Apress, 2008, $54.99, ISBN 1-59059-955-1. ACM Queue: Tomorrow’s Computing Today, 6 (5):59, September 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Nagel:2014:CGE

[NBV14] Fabian Nagel, Gavin Bierman, and Stratis D. Viglas. Code generation for efficient query processing in managed runtimes. Proceedings of the VLDB Endowment, 7(12):1095–1106, August 2014. CODEN ???? ISSN 2150-8097.

Nagel:2006:PC

[NEG^{+}06] Christian Nagel, Bill Evjen, Jay Glynn, Karli Watson, Morgan Skinner, and Allen Jones, editors. Professional C# 2005. John Wiley, New York, NY, USA, 2006. ISBN 0-7645-7534-1. xlvi + 1540 pp. LCCN QA76.73.C154 P82 2005. URL http://www.wiley.com/

WileyCDA/WileyTitle/productCd-0764575341.html.

Nagel:2010:PCN

[NEG^{+}10] Christian Nagel, Bill Evjen, Jay Glynn, Karli Watson, and Morgan Skinner, editors. Professional C# 4 and .NET 4. Wrox programmer to programmer. John Wiley, New York, NY, USA, 2010. ISBN 0-470-50225-8. lix + 1474 pp. LCCN QA76.73.C154.

Newman:2005:S

[New05] Chris Newman. SQLite. Developer’s library. Howard W. Sams, Indianapolis, IN 46268, USA, 2005. ISBN 0-672-32685-X (paperback), 0-7686-6366-0 (e-book). xvi + 313 pp. LCCN QA76.73.S67

N488 2005eb. URL http://proquest.safaribooksonline.com/
067232685X;   http://proquestcombo.safaribooksonline.com/
067232685X;   http://www.contentreserve.com/TitleInfo.asp?
ID={5D380338-820F-4EC3-BAA3-668B6BF49170}\%26Format=50;
http://www.myilibrary.com?id=108361.

Nilsson:2006:ADD

[Nil06] Jimmy Nilsson. Applying domain-driven design and patterns with examples in C# and .NET. Addison-Wesley, Reading, MA, USA, 2006. ISBN 0-321-26820-2. ???? pp. LCCN QA76.76.D47 N645 2006. URL

http://www.loc.gov/catdir/toc/ecip068/2006004371.html.

Nodelman:2004:OCC

[Nod04] Vladimir Nodelman. OOP via C++, C#, …? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (3):255, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Onion:2003:EANb

[Oni03a] Fritz Onion. Essential ASP.NET with examples in C#. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-76040-1. xxiv + 393 pp. LCCN TK5105.8885.A26 O53 2003.

Onion:2003:EANa

[Oni03b] Fritz Onion. Essential ASP.NET with examples in Visual Basic .NET. Microsoft .NET development series. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-76039-8. xxiv + 395 pp. LCCN TK5105.8885.A26 O55 2003.

Pelland:2005:MVC

[Pel05] Patrice Pelland. Microsoft Visual C# 2005 Express Edition: build a program now! Microsoft Press, Redmond, WA, USA, 2005. ISBN 0-7356-2229-9. ???? pp. LCCN ????

Pelland:2006:MVC

[Pel06a] Patrice Pelland. Microsoft Visual C# 2005: express edition. Academic Service, Den Haag, The Netherlands, 2006. ISBN 90-395-2439-4. xiii + 178 pp. LCCN ????

Pelland:2006:PJM

[Pel06b] Patrice Pelland. Programmieren Sie jetzt! — Microsoft Visual C# 2005 Express Edition. Programmieren Sie jetzt! Microsoft Press, Unterschleicheim, Germany, 2006. ISBN 3-86063-569-7. xi + 211 pp. LCCN ???? EUR 19.90 (DE). URL http://deposit.ddb.de/

cgi-bin/dokserv?id=2669514&prov=M&dok_var=1&dok_ext=htm.

Penton:2005:BCG

[Pen05] Ron Penton. Beginning C# game programming. Course Technology PTR, Boston, MA, USA, 2005. ISBN 1-59200-517-9. xix + 318 pp.

LCCN QA76.73.C15 P46 2005eb. URL http://www.columbia.edu/

cgi-bin/cul/resolve?clio5389454.

Perry:2006:CCN

[Per06] Stephen C. Perry. Core C# and .NET. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2006. ISBN 0-13-147227-5. xxx + 972 pp. LCCN QA76.73.C153 P468 2005. URL http://www.loc.

gov/catdir/toc/ecip0516/2005021301.html.

Petzold:2002:PMW

[Pet02] Charles Petzold. Programming Microsoft Windows with C#. Microsoft Press, Redmond, WA, USA, 2002. ISBN 0-7356-1370-2. xxii + 1290 pp. LCCN QA76.73.C154 P48 2002.

Perelman:2012:TDC

[PGBG12] Daniel Perelman, Sumit Gulwani, Thomas Ball, and Dan Grossman. Type-directed completion of partial expressions. ACM SIGPLAN Notices, 47(6):275–286, June 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PLDI ’12 proceedings.

Pattison:2003:BAC

[PH03] Ted Pattison and Joe Hummel. Building applications and components with Visual Basic .NET. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-73495-8. xxviii + 562 pp. LCCN

QA76.73.B3 P257 2003. URL http://www.loc.gov/catdir/toc/
ecip047/2003016759.html.

Pizlo:2008:SCR

[PPS08] Filip Pizlo, Erez Petrank, and Bjarne Steensgaard. A study of concurrent real-time garbage collectors. ACM SIGPLAN Notices, 43 (6):33–44, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Prabhu:2010:SPS

[PRV10] Prakash Prabhu, Ganesan Ramalingam, and Kapil Vaswani. Safe programmable speculative parallelism. ACM SIGPLAN Notices, 45 (6):50–61, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Perez:2017:UCP

[PS17] Alejandro Pérez and Pablo Sánchez. On the use of C# partial classes for the implementation of software product lines. The Computer Journal, 60(1):86–109, January 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Petzold:2002:MWP

[PSN^{+}02] Charles Petzold, Henk Scholten, Wout Natter, et al. Microsoft Windows programmeren met C#. (Dutch) [Microsoft Windows programming with C#]. Academic Service, Schoonhoven, The Netherlands, 2002. ISBN 90-395-1940-4. xiv + 1312 pp. LCCN ????

Pinheiro:2020:MCA

[PVR^{+}20] Pedro Pinheiro, José Carlos Viana, Márcio Ribeiro, Leo Fernandes, Fabiano Ferrari, Rohit Gheyi, and Baldoino Fonseca. Mutating code annotations: an empirical evaluation on Java and C programs. Science of Computer Programming, 191(??):??, June 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0167642320300290.

Powell:2001:CNF

[PW01] Robert Powell and Richard Weeks. C# and the .NET framework: the C++ perspective. Howard W. Sams, Indianapolis, IN 46268, USA, 2001. ISBN 0-672-32153-X (paperback). xi + 620 pp. LCCN QA76.73.C154 P69 2002; QA76.73.C154 .P63 2001.

Pizlo:2010:SFT

[PZM^{+}10] Filip Pizlo, Lukasz Ziarek, Petr Maj, Antony L. Hosking, Ethan Blanton, and Jan Vitek. SCHISM: fragmentation-tolerant real-time garbage collection. ACM SIGPLAN Notices, 45(6):146–159, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Robinson:2002:CVP

[R^{+}02] Simon Robinson et al., editors. C# voor professionals. (Dutch) [C# for Professionals]. Academic Service, Schoonhoven, The Netherlands, 2002. ISBN 90-395-1962-5. xix + 1258 pp. LCCN ????

Robinson:2003:PCS

[RAC^{+}03] Simon Robinson, K. Scott Allen, Ollie Cornes, Jay Glynn, Zach Greenvoss, Burton Harvey, Christian Nagel, Morgan Skinner, Karli Watson, and Steve Danielson, editors. Professional C#. Programmer to programmer. John Wiley, New York, NY, USA, second edition, 2003. ISBN 0-7645-4398-9. xxiii + 1223 pp. LCCN QA76.73.C154 P76 2002b.

Reges:2002:CCR

[Reg02a] S. Reges. Can C# replace Java in CS1 and CS2? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (3):4–8, 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reges:2002:SFI

[Reg02b] Stuart Reges. SIGCSE 2002 forum: industry in the ivory tower. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):367, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Reilly:2002:ANS

[Rei02] Douglas Reilly. ASP.NET server components. Dr. Dobb’s Journal of Software Tools, 27(6):62, 64, 66–68, June 2002. CODEN

DDJOEB. ISSN 1044-789X. URL http://www.ddj.com/ftp/2002/
2002_06/aspnet.txt;   http://www.ddj.com/ftp/2002/2002_
06/aspnet.zip.

Reynolds-Haertle:2002:POM

[RHSvS02] Robin A. Reynolds-Haertle, Maarten Stefanski, and Marijke van Steenis. Programmeercursus OOP met MS VB.NET en MS Visual C#. (Dutch) [Object-oriented programming course with Microsoft VB.NET and Microsoft Visual C#. Academic Service, Schoonhoven, The Netherlands, 2002. ISBN 90-395-1960-9. xv + 420 pp. LCCN ????

Richter:2006:CCS

[Ric06] Jeffrey Richter. CLR via C#. Microsoft Press, Redmond, WA, USA, second edition, 2006. ISBN 0-7356-2163-2. xxviii + 693 pp. LCCN ???? Edited by Devon Musgrave.

Richter:2010:CCH

[Ric10] Jeffrey Richter. CLR via C#. Microsoft Press, Redmond, WA, USA, third edition, 2010. ISBN 0-7356-4046-7, 0-7356-2704-5. xviii + 873 pp. LCCN ????

Rauniyar:2017:DDR

[RIS^{+}17] Ashish Rauniyar, Mohammad Irfan, Oka Danil Saputra, Jin Woo Kim, Ah Ra Lee, Jae Min Jang, and Soo Young Shin. Design and development of a real-time monitoring system for multiple lead-acid batteries based on Internet of Things. Future Internet, 9(3):28, June 29, 2017. CODEN ???? ISSN 1999-5903. URL https://www.mdpi.

com/1999-5903/9/3/28.

Robinson:2004:PCS

[RNG^{+}04] Simon Robinson, Christian Nagel, Jay Glynn, Morgan Skinner, Karli Watson, and Bill Evjen, editors. Professional C#. Programmer to programmer. John Wiley, New York, NY, USA, third edition, 2004. ISBN 0-7645-5759-9. xxxvii + 1356 pp. LCCN QA76.73.C154P76 2004.

Robinson:2002:PCS

[Rob02] Simon Robinson, editor. Professional C#. Programmer to programmer. Wrox Press, Chicago, IL, USA, second edition, 2002. ISBN 1-86100-704-3. xxiii + 1223 pp. LCCN ????

Rajan:2010:GPM

[RRY10] Kaushik Rajan, Sriram Rajamani, and Shashank Yaduvanshi. GUESSTIMATE: a programming model for collaborative distributed systems. ACM SIGPLAN Notices, 45(6):210–220, June 2010. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Russo:2008:JPV

[Rus08] Claudio V. Russo. Join patterns for Visual Basic. ACM SIGPLAN Notices, 43(10):53–72, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ruvalcaba:2004:BYO

[Ruv04] Zak Ruvalcaba. Build your own ASP.NET Website using C# and VB.NET. SitePoint Pty. Ltd., Collingwood, Victoria, Australia, 2004. ISBN 0-9579218-6-1. xviii + 746 pp. LCCN TK5105.8885.A26 R89 2004b. URL http://www.columbia.edu/

cgi-bin/cul/resolve?clio5459110.

Ramalingam:2013:FTI

[RV13] Ganesan Ramalingam and Kapil Vaswani. Fault tolerance via idempotence. ACM SIGPLAN Notices, 48(1):249–262, January 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reynolds:2002:NED

[RW02] Matthew Reynolds and Karli Watson. .NET enterprise development with C# from design to deployment. Programmer to programmer. Wrox Press, Chicago, IL, USA, 2002. ISBN 1-86100-591-1. ix + 453 pp. LCCN QA76.76.M52 R496 2002b.

Saxena:2009:PEO

[SA09] Vipin Saxena and Deepak Arora. Performance evaluation for object oriented software systems. ACM SIGSOFT Software Engineering Notes, 34(2):1–5, March 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saxon:2003:OOR

[Sax03] Charles S. Saxon. Object-oriented recursive descent parsing in C#. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):82–85, December 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/2003.bib.

Shankar:2007:DAI

[SB07] Ajeet Shankar and Rastislav Bodík. DITTO: automatic incrementalization of data structure invariant checks (in Java). ACM SIG-PLAN Notices, 42(6):310–319, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schildt:2001:CBG

[Sch01] Herbert Schildt. C#: a beginner’s guide. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-213329-5. xviii + 621 pp. LCCN QA76.73.C154 S338 2001. URL http://osborne.

com/programming\%5Fwebdev/0072133295/0072133295.shtml;
http://www.loc.gov/catdir/description/mh024/2002277354.
html;   http://www.loc.gov/catdir/toc/mh023/2002277354.
html.

Schildt:2006:CCR

[Sch06a] Herbert Schildt. C# 2.0: the complete reference. Osborne/McGraw-Hill, Berkeley, CA, USA, second edition, 2006. ISBN 0-07-226209-5 (paperback), 0-07-148373-X (electronic). xxi + 890 pp. LCCN QA76.73.C154 S35 2006.

Schumann:2006:VCK

[Sch06b] Hans-Georg Schumann. Visual C# 2005 für Kids. bhv, Heidelberg, Germany, 2006. ISBN 3-8266-8626-8. 361 pp. LCCN ???? EUR 24.95.

Schildt:2010:CCR

[Sch10] Herbert Schildt. C# 4.0: the complete reference. McGraw-Hill, New York, NY, USA, 2010. ISBN 0-07-174116-X (paperback). xxv + 949 pp. LCCN QA76.73.C154. US$34.99.

Sphar:2008:CD

[SD08] Chuck Sphar and Stephen R. Davis. C# 2008 for dummies. For dummies. John Wiley, New York, NY, USA, 2008. ISBN 0-470-19109-0 (paperback). xx + 434 pp. LCCN

QA76.73.C154 S68 2008. URL http://www.loc.gov/catdir/
enhancements/fy0808/2008920770-b.html;   http://www.loc.
gov/catdir/enhancements/fy0808/2008920770-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0810/2008920770-t.html.

Sallenave:2012:LGE

[SD12] Olivier Sallenave and Roland Ducournau. Lightweight generics in embedded systems through static analysis. ACM SIGPLAN Notices, 47(5):11–20, May 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCTES ’12 proceedings.

Sells:2004:WFPa

[Sel04] Chris Sells. Windows forms programming in C#. Microsoft .NET development series. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-321-11620-8. xliii + 681 pp. LCCN QA76.76.M52 S45 2004.

Sestoft:2008:PLC

[Ses08] Peter Sestoft. Programming language concepts for software developers. ACM SIGPLAN Notices, 43(11):119–123, November 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stallings:2001:PBB

[SG01] William Stallings and Lou Grinzo. Programmer’s bookshelf: a Bluetooth trio and C#. Dr. Dobb’s Journal of Software Tools, 26(9):130, 132, September 2001. CODEN DDJOEB. ISSN 1044-789X. URL

http://www.ddj.com/.

Sells:2004:WFPb

[SG04] Chris Sells and Justin Gehtland. Windows forms programming in Visual Basic .NET. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-321-12519-3. xlvi + 680 pp. LCCN QA76.76.M52 S46 2003.

Stellman:2008:HFC

[SG08] Andrew Stellman and Jennifer Greene. Head first C#. O’Reilly’s Head first series. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51482-4 (paperback). xl + 738 pp. LCCN QA76.73.C154; QA76.73.C154eb.

URL http://www.oreilly.com/catalog/9780596514822.

Stellman:2010:HFC

[SG10] Andrew Stellman and Jennifer Greene. Head first C#. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2010. ISBN 1-4493-8034-4 (paperback). ???? pp. LCCN ???? US$35.50.

Sestoft:2004:CP

[SH04] Peter Sestoft and Henrik I. Hansen. C# precisely. MIT Press, Cambridge, MA, USA, 2004. ISBN 0-262-69317-8 (paperback). viii + 204 pp. LCCN QA76.73.C154.S47 2004; QA76.73.C154.S47.

Sharp:2006:MVCb

[Sha06a] John Sharp. Microsoft Visual C# 2005: étape par étape. (French) [Microsoft Visual C# 2005: Step by step]. Microsoft Press, Les Ulis, France, 2006. ISBN 2-10-049606-9. x + 522 pp. LCCN ???? Translated by Chantal Kolb, Dorothée Sittler, and Christine Eberhardt.

Sharp:2006:MVCa

[Sha06b] John Sharp. Microsoft Visual C# 2005 step by step. Microsoft Press, Redmond, WA, USA, 2005 edition, 2006. ISBN 0-7356-2129-2 (paperback). xxvii + 555 pp. LCCN QA76.73.C154 S53 2006.

Shu:2005:FCD

[Shu05] Ge Shu. F# and C# on the .NET platform. Master’s thesis, Department of Computer Science, University of Wales, Swansea, UK, 2005.

Sharp:2002:PMV

[SJNH02] John Sharp, John Jagger, Feico Nater, and Wim Huberts. Programmeercursus Microsoft Visual C# .NET. (Dutch) [Microsoft Visual C# .NET programming course]. Programmeercursus; Programmeercursus. Academic Service, Schoonhoven, The Netherlands, 2002. ISBN 90-395-1961-7. xxvi + 611 pp. LCCN ????

Sauer:2008:VCH

[SK08] Florian Sauer and Joachim Karl, editors. Visual C# 2008. Addison-Wesley, München, Germany, 2008. ISBN 3-8273-6114-1. ???? pp. LCCN ????

Santhiar:2017:SDD

[SK17] Anirudh Santhiar and Aditya Kanade. Static deadlock detection for asynchronous C# programs. ACM SIGPLAN Notices, 52(6): 292–305, June 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Skeet:2010:CD

[Ske10] Jon Skeet. C# in Depth. Manning Publications, Greenwich, CT, USA, 2010. ISBN 1-935182-47-1. xxx + 554 pp. LCCN QA76.73.C154 S58 2011. US$49.99.

Stubblebine:2008:RAK

[SKS08] Toni Stubblebine, Peter Klicman, and Lars Schulten. Reguläre Ausdrücke — kurz and gut [für Perl, Ruby, PHP, C#, Python, Java and .NET]. O’Reillys Taschenbibliothek. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2008. ISBN 3-89721-535-7. 131 pp. LCCN ???? EUR 9.90.

Spring:2010:RAI

[SPP^{+}10] Jesper Honig Spring, Filip Pizlo, Jean Privat, Rachid Guerraoui, and Jan Vitek. Reflexes: Abstractions for integrating highly responsive tasks into Java applications. ACM Transactions on Embedded Computing Systems, 10(1):4:1–4:??, August 2010. CODEN ???? ISSN 1539-9087.

Singh:2013:TGC

[SSK13] Pavitdeep Singh, Satwinder Singh, and Jatinder Kaur. Tool for generating code metrics for C# source code using abstract syntax tree technique. ACM SIGSOFT Software Engineering Notes, 38(5):1–6, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sorensen:2002:AND

[SSRW02] Ronan Sorensen, George Shepherd, John Roberts, and Russ Williams, editors. Applied .NET: developing people-oriented software using C#. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-73828-7 (paperback). xviii + 397 pp. LCCN QA76.73.C154 A67 2002. US$30.99.

Stark:2005:FSV

[Stä05] Robert F. Stärk. Formal specification and verification of the C# thread model. Theoretical Computer Science, 343(3):482–508, October 17, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stevens:2002:CPCb

[Ste02a] Al Stevens. C programming: C# and the outhouse paradigm. Dr. Dobb’s Journal of Software Tools, 27(7):79–81, July 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.ddj.com/.

Stevens:2002:CPW

[Ste02b] Al Stevens. C programming: Wave repairs and a first look at C#. Dr. Dobb’s Journal of Software Tools, 27(3):77–78, 80, March 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.ddj.com/

ftp/2002/2002_03/cpro0302.txt;   http://www.ddj.com/ftp/
2002/2002_03/cpro0302.zip.

Siek:2015:BCT

[STW15] Jeremy Siek, Peter Thiemann, and Philip Wadler. Blame and coercion: together again for the first time. ACM SIGPLAN Notices, 50 (6):425–435, June 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Subramaniam:2005:NGW

[Sub05] Venkat Subramaniam. .NET gotchas: 75 ways to improve your C# and VB.NET programs. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00909-7. xviii + 372 pp. LCCN QA76.625 .S93 2005.

Shacham:2009:CAS

[SVY09] Ohad Shacham, Martin Vechev, and Eran Yahav. Chameleon: adaptive selection of collections. ACM SIGPLAN Notices, 44(6):408–418, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Swaine:2001:PPA

[Swa01] Michael Swaine. Programming paradigms: Alternatives to Java. Dr. Dobb’s Journal of Software Tools, 26(2):135–137, February 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.ddj.com/.

Sweeney:2006:NMP

[Swe06] Tim Sweeney. The next mainstream programming language: a game developer’s perspective. ACM SIGPLAN Notices, 41(1):269, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Swamy:2013:VHO

[SWS^{+}13] Nikhil Swamy, Joel Weinberger, Cole Schlesinger, Juan Chen, and Benjamin Livshits. Verifying higher-order programs with the Dijkstra monad. ACM SIGPLAN Notices, 48(6):387–398, June 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tagliaferri:2003:CC

[Tag03] Matt Tagliaferri, editor. C# complete. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2003. ISBN 0-7821-4203-6 (paperback). xxvii + 974 pp. LCCN QA76.73.C154

C454 2003b. US$14.99. URL http://www.loc.gov/catdir/
enhancements/fy0613/2002115481-b.html.

Tate:2010:IOO

[TCH10] Ross Tate, Juan Chen, and Chris Hawblitzel. Inferable objectoriented typed assembly language. ACM SIGPLAN Notices, 45(6): 424–435, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Telles:2002:CBB

[Tel02] Matthew A. Telles. C# black book. Coriolis Group Books, Scottsdale, AZ, USA, 2002. ISBN 1-58880-192-6 (paperback/CD-ROM). xxii + 643 pp. LCCN QA76.73.C154 T45 2002eb. URL http://www.

columbia.edu/cgi-bin/cul/resolve?clio5390553.

Telles:2006:VCB

[Tel06] Matthew Telles. Visual C# 2005 Black Book. Pagina Förlags AB, Box 706, Travgatan 92, 194 27 Upplands-Väsby, Sweden, 2006. ISBN 1-933097-16-7. 680 pp. LCCN ????

Titus:2003:CTH

[TFRS03] Tobin Titus, Fabio Claudio Ferracchiati, Tejaswi Redkar, and Srinivasa Sivakumar, editors. C# threading handbook. Programmer to programmer. Wrox Press, Chicago, IL, USA, 2003. ISBN 1-86100-829-5. v + 272 pp. LCCN QA76.73.C154 T58 2004b.

Teilhet:2004:CC

[TH04] Stephen Teilhet and Jay Hilyard. C# cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00339-0. xxii + 831 pp. LCCN QA76.73.C154 T43 2004.

Telles:2002:C

[THG^{+}02] Matthew Telles, Dré Holthuijsen, Gert Gielen, et al. C#. Easy Computing, Brussels, Belgium, 2002. ISBN 90-5167-459-7. 607 pp. LCCN ????

Thomsen:2002:DPC

[Tho02] Carsten Thomsen. Database programming with C#. Books for professionals by professionals. Apress, Berkeley, CA, USA, 2002. ISBN 1-59059-010-4. xix + 667 pp. LCCN QA76.73.C154 T46 2002b.

Titzer:2013:HCF

[Tit13] Ben L. Titzer. Harmonizing classes, functions, tuples, and type parameters in Virgil III. ACM SIGPLAN Notices, 48(6):85–94, June 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Troelsen:2001:CNP

[Tro01] Andrew W. Troelsen. C# and the .NET platform. Apress, Berkeley, CA, USA, 2001. ISBN 1-893115-59-3. xxiii + 970 pp. LCCN

QA76.73.C154 T764 2001b. URL http://www.columbia.edu/cgi-
bin/cul/resolve?clio3333214.

Troelsen:2003:CNP

[Tro03] Andrew W. Troelsen. C# and the .NET platform. Apress, Berkeley, CA, USA, second edition, 2003. ISBN 1-59059-055-4. 1158 (est.) pp.

LCCN QA76.73.C154 T764 2003b. URL http://www.columbia.

edu/cgi-bin/cul/resolve?clio4905456.

Troelsen:2005:PCN

[Tro05] Andrew W. Troelsen. Pro C# 2005 and the .NET 2.0 platform. Apress, Berkeley, CA, USA, third edition, 2005. ISBN 1-59059-419-3. xlv + 982 pp. LCCN QA76.73.C153 T764 2005.

Troelsen:2010:PCN

[Tro10] Andrew Troelsen. Pro C# 2010 and the .Net 4.0 platform. Apress, Berkeley, CA, USA, fifth edition, 2010. ISBN 1-4302-2550-5, 1-4302-2549-1. xxxviii + 1712 pp. LCCN QA76.73.C154 T764 2010.

Tsafrir:2009:MDW

[TWBS09] Dan Tsafrir, Robert W. Wisniewski, David F. Bacon, and Bjarne Stroustrup. Minimizing dependencies within generic classes for faster and smaller programs. ACM SIGPLAN Notices, 44(10):425–444, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ullman:2004:BAN

[UKH^{+}04] Chris Ullman, John Kauffman, Chris Hart, Dave Sussman, and Dan Maharry, editors. Beginning ASP.NET 1.1 with Visual C# .NET 2003. John Wiley, New York, NY, USA, 2004. ISBN 0-7645-5708-4. ???? pp. LCCN TK5105.8885.A26 B452 2004b. URL http://www.

columbia.edu/cgi-bin/cul/resolve?clio4905523.

Vaughn:2002:ANE

[VB02] William R. Vaughn and Peter Blackburn. ADO.NET examples and best practices for C# programmers. Apress, Berkeley, CA, USA, 2002. ISBN 1-59059-012-0. ???? pp. LCCN QA76.625 .V384

2002b. URL http://www.columbia.edu/cgi-bin/cul/resolve?
clio4905480.

Vick:2004:VBN

[Vic04] Paul Vick. The Visual Basic .NET programming language. Microsoft .NET development series. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-321-16951-4. xxiv + 407 pp. LCCN QA76.73.B3 V484 2004.

Watson:2005:BVC

[W^{+}05] Karli Watson et al., editors. Beginning Visual C# 2005. John Wiley, New York, NY, USA, 2005. ISBN (paper/website), 0-7645-7847-2 (paper/website). ???? pp. LCCN QA76.73.C154 B483

2005. URL http://www.loc.gov/catdir/enhancements/fy0622/
2005010692-d.html;   http://www.loc.gov/catdir/toc/ecip0511/
2005010692.html.

Watson:2006:BVC

[W^{+}06] Karli Watson et al., editors. Beginning Visual C# 2005. Wrox beginning guides. John Wiley, New York, NY, USA, 2006. ISBN 0-7645-7847-2 (paper/website). xxxix + 1062 pp. LCCN QA76.73.C154

B483 2006eb; **See. URL http://link.library.utoronto.ca/
eir/EIRdetail.cfm?Resources\_\_ID=369413&T=; http://www.
loc.gov/catdir/toc/ecip0511/2005010692.html; resource.

Wagner:2002:CCL

[Wag02] Bill Wagner. C# core language: little black book. Little black book. Coriolis Group Books, Scottsdale, AZ, USA, 2002. ISBN 1-58880-058-X. xxvii + 446 pp. LCCN QA76.73.C154 W34 2002eb. URL

http://www.columbia.edu/cgi-bin/cul/resolve?clio5375250.

Wagner:2005:ECS

[Wag05] Bill Wagner. Effective C#: 50 specific ways to improve your C#. Effective software development series. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-321-24566-0. xviii + 307 pp. LCCN QA76.73.C154 W343 2005.

Wagner:2010:ECH

[Wag10] Bill Wagner. Effective C#: 50 specific ways to improve your C#. Effective software development series. Addison-Wesley, Reading, MA, USA, second edition, 2010. ISBN 0-321-65914-7, 0-321-65870-1. xviii + 328 pp. LCCN ????

Watkins:2003:PNE

[WAH03] Damien Watkins, Brad Abrams, and Mark (Mark J.) Hammond. Programming in the .NET environment. Microsoft .NET development series. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-77018-0 (paperback). xxviii + 523 pp. LCCN QA76.625 .W38 2002.

Wang:2011:EEU

[Wan11] Alf Inge Wang. Extensive evaluation of using a game project in a software architecture course. ACM Transactions on Computing Education, 11(1):5:1–5:??, February 2011. CODEN ???? ISSN 1946-

Watson:2006:BCD

[Wat06] Karli Watson. Beginning C# 2005 Databases. Wrox beginning guides. Wiley/Wrox, Indianapolis, IN, USA, 2006. ISBN 0-470-04406-3 (paperback). xxiii + 501 pp. LCCN QA76.73.C154 W39

2006. URL http://www.loc.gov/catdir/enhancements/fy0741/
2006025170-b.html;   http://www.loc.gov/catdir/enhancements/
fy0741/2006025170-d.html;   http://www.loc.gov/catdir/toc/
ecip0618/2006025170.html.

Watson:2010:CH

[Wat10a] Ben Watson. C# 4.0 how-to: Real Solutions for C# 4.0 programs. Howard W. Sams, Indianapolis, IN 46268, USA, 2010. ISBN 0-672-33063-6 (paperback). xiii + 653 pp. LCCN QA76.73.C154 W38 2010.

Watson:2010:BMV

[Wat10b] Karli Jane Watson. Beginning Microsoft Visual C# 2010. John Wiley, New York, NY, USA, 2010. ISBN 0-470-50226-6 (paperback). ???? pp. LCCN ????

Wille:2001:CI

[WdC01] Christoph Wille and Sebastiaan de Coo. C#: een inleiding. (Dutch) [C#: an introduction]. SAMS, Amsterdam, The Netherlands, 2001. ISBN 90-430-0425-1. x + 204 pp. LCCN ????

Watson:2002:BVC

[WEG^{+}02] Karli Watson, David Espinosa, Zach Greenvoss, Jacob Hammer Pedersen, Jon Reid, Christian Nagel, Matthew Reynolds, Morgan Skinner, and Eric White, editors. Beginning Visual C#. Programmer to programmer. Wrox Press, Chicago, IL, USA, revised edition, 2002. ISBN 1-86100-758-2, 0-7645-4382-2. xvii + 903 pp. LCCN QA76.73.C15 B44 2002b.

Wiltamuth:2000:CPL

[Wil00] Scott Wiltamuth. The C# programming language. Dr. Dobb’s Journal of Software Tools, 25(10):21–22, 24, 26, 28, 30–32, October 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.ddj.com/

ftp/2000/2000_10/cshrp.txt.

Wilson:2001:PBT

[Wil01] Gregory V. Wilson. Programmer’s bookshelf: Time warps. Dr. Dobb’s Journal of Software Tools, 26(2):159–160, February 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.ddj.com/.

Wildermuth:2003:PAN

[Wil03] Shawn Wildermuth. Pragmatic ADO.NET: data access for the Internet world. Microsoft .NET development series. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-74568-2 (paperback). xvii + 357 pp. LCCN QA76.9.D26 W55 2003.

Winkler:2004:CCJ

[Win04] J. F. H. Winkler. C#: a competitor of Java? It, 46(2):85–93, 2004. CODEN ???? ISSN 1611-2776.

Willcock:2005:UMC

[WLR05] Jeremiah Willcock, Andrew Lumsdaine, and Arch Robison. Using MPI with C# and the Common Language Infrastructure. Concurrency and Computation: Practice and Experience, 17(7–8):895–917, June/July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Willms:2003:CZS

[WNG03] Roland Willms, Jan J. Nathan, and Gert Gielen. C# zonder stress. (Dutch) [C# without stress]. Easy Computing, Brussels, Belgium, 2003. ISBN 90-5167-417-1. 268 pp. LCCN ????

Worner:2004:ACF

[Wor04] Martin Worner. Applied C# in financial markets. Wiley finance series. John Wiley, New York, NY, USA, 2004. ISBN 0-470-87061-3. xviii + 119 pp. LCCN QA76.73.C15 W67 2004. US$45.00.

Wehr:2010:JBP

[WT10] Stefan Wehr and Peter Thiemann. JavaGI in the battlefield: practical experience with generalized interfaces. ACM SIGPLAN Notices, 45(2):65–74, February 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:2009:SCC

[XJC09] Dana N. Xu, Simon Peyton Jones, and Koen Claessen. Static contract checking for Haskell. ACM SIGPLAN Notices, 44(1):41–52, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yao:2004:NCF

[YD04] Paul Yao and David Durant. .NET Compact Framework programming with C#. Microsoft .NET development series. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-321-17403-8. xl + 1379 pp. LCCN

QA76.73.C154 Y36 2004. URL http://www.loc.gov/catdir/toc/
ecip0414/2004002429.html.

Yang:2010:SLI

[YH10] Jean Yang and Chris Hawblitzel. Safe to the last instruction: automated verification of a type-safe operating system. ACM SIGPLAN Notices, 45(6):99–110, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zyulkyarov:2010:DPU

[ZHU^{+}10] Ferad Zyulkyarov, Tim Harris, Osman S. Unsal, Adrían Cristal, and Mateo Valero. Debugging programs that use atomic blocks and transactional memory. ACM SIGPLAN Notices, 45(5):57–66, May 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhao:2014:CSP

[ZWZ^{+}14] Zhijia Zhao, Bo Wu, Mingzhou Zhou, Yufei Ding, Jianhua Sun, Xipeng Shen, and Youfeng Wu. Call sequence prediction through probabilistic calling automata. ACM SIGPLAN Notices, 49(10): 745–762, October 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).